### v1

# JWST Discovery of a High-Redshift Tidal Disruption Event Candidate in COSMOS-Web

Abstract
The rates and properties of tidal disruption events (TDEs) provide valuable insights into their host galaxy central stellar densities and the demographics of their central supermassive black holes (SMBHs). TDEs have been observed only at low redshifts ($z\lesssim 1$), due to the difficulty in conducting deep time-domain surveys. In this work, we present the discovery of a high-redshift TDE candidate, HZTDE-1, in the COSMOS-Web survey with JWST’s NIRCam, using a novel selection technique based on color and morphology. We first outline a methodology for identifying high-z TDEs in deep infrared imaging surveys, leveraging their unique spectral energy distributions (SEDs) and morphologies of these transients. We apply this technique to COSMOS-Web in filters F115W, F150W, F277W, and F [^4] W, and identify HZTDE-1, a transient point source relative to archival UltraVISTA infrared observations. If we assume it is a TDE, we estimate its photometric redshift to be $z=5.02^{+1.32}_{-1.11}$. HZTDE-1 cannot be explained by reasonable supernova or AGN models. However, we cannot rule out a superluminous supernova at $z\gtrsim 3$. If confirmed with follow-up observations, HZTDE-1 would represent the highest-redshift TDE discovery to date, and would suggest an enhancement of the TDE rate in the high-redshift universe. Our method, which can be applied to future deep surveys with JWST and Roman, offers a pathway to identify TDEs at $z>4$ and probe black hole demographics at early cosmic times.
Supermassive black holes, Tidal disruption, Infrared surveys, James Webb Space Telescope
1 Introduction
A tidal disruption event (TDE) occurs when a star passes close enough to a supermassive black hole (SMBH) that the tidal forces on the star overcome its self-gravity, and destroy the star (Hills, 1975; Rees, 1988; Evans & Kochanek, 1989; Ulmer, 1999). This process leads to a luminous flare ($\sim 10^{43}$erg/s), as approximately half of the mass of the star is accreted and the other half is ejected at high velocity (Strubbe & Quataert, 2009; Gezari et al., 2012). This highly transient accretion event contrasts with the more long term, sustained accretion that is associated with active galactic nuclei (AGN) (Begelman et al., 1984; Krolik, 1999; Event Horizon Telescope Collaboration et al., 2019). Despite decades of dedicated efforts, several outstanding questions surrounding the formation and evolution of SMBHs persist. In particular, the uncertain origin of the first SMBHs has outlined the need for robustly measuring the early Universe SMBH mass distribution (Volonteri et al., 2003; Begelman et al., 2006). TDEs are one of the only phenomena, alongside AGN, that present a luminous observable very close to a SMBH. Unlike AGN, which prefer very massive SMBH hosts (Kauffmann et al., 2003), TDEs of main-sequence stars by non-rotating (Schwarzschild) black holes can only occur for lower mass SMBHs because the disrupting black hole has an upper mass limit of $M_{BH}<10^{8}M_{\odot}$ (Hills, 1975). Beyond this mass, the tidal radius $R_{t}=R_{\star}(M_{BH}/M_{\star})$ is within the Schwarzschild radius $R_{s}=2GM/c^{2}$. The TDE rate’s sensitivity to the black hole mass make TDEs an excellent probe of the SMBH mass function (Stone & Metzger, 2016; Kochanek, 2016). The TDE rate is also affected by the stellar mass function, so a redshift-evolving TDE rate could also be sensitive to the initial mass function (IMF) (Magorrian & Tremaine, 1999). Additionally, stellar dynamics calculations imply that galaxies with lower-mass SMBHs have higher intrinsic TDE rates (Stone & Metzger, 2016). Therefore, TDEs are ideal probes of distant, lower-mass SMBHs and the stars which surround them.
The primary tool used to measure the masses of high-redshift SMBHs, when dynamical information is not available, is spectroscopy of AGN (Peterson, 1998; Kaspi et al., 2000). Due to Malmquist bias, the most massive tail of the high redshift BH mass function is observed, leading to observations of SMBHs which are overmassive relative to their host galaxies (Pacucci & Loeb, 2024) and their cosmic epoch (Juodžbalis et al., 2024). JWST observations are able to sample lower-luminosity AGN and thus lower-mass BHs (e.g. Li et al., 2025; Harikane et al., 2023), but are prone to systematic biases in virial mass estimation methods (Krolik, 2001; Shen, 2013; Bertemes et al., 2025). TDEs offer an independent tool to study the less massive end of this mass function. The peak luminosities of TDEs (Mockler et al., 2019), their fallback rates (Bandopadhyay et al., 2024), and their late-time plateau luminosities (Mummery et al., 2024) have all been linked to their host SMBH mass. Furthermore, TDE rates can be a strong probe of the fraction of black holes in low-mass galaxies (Stone & Metzger, 2016). TDEs are transients, which makes them significantly more difficult to observe. While thousands of AGN have been observed up to $z=11$ (Maiolino et al., 2023), but only of order 100 TDEs have been spectroscopically confirmed at the time of writing this paper (Gezari, 2021), with the highest redshift TDE the jetted TDE 2022cmc at $z=1.19$ (Andreoni et al., 2022).
Typically TDEs are discovered through wide-field, time-domain optical surveys (e.g. the Zwicky Transient Facility (ZTF) Graham et al., 2019; Bellm et al., 2019). They are slow (lasting 100100100100200200200200 days), blue (peaking in the UV) transients that do not typically cool as they fade (Gezari, 2021). TDEs last on scales of months, typically fading from peak in $\sim 200$ days (van Velzen et al., 2021). The origin of the early-time UV/optical emission is debated, and still not fully understood (Ryu et al., 2023; Dai et al., 2018; Roth et al., 2020). In this work, we adopt a fully empirical, observation-driven model of TDE spectra and light curves.
At higher redshifts, TDEs are fainter and therefore their light curves are more difficult to observe far from maximum brightness. They are time-dilated and therefore evolve very slowly at higher redshifts. At $z>4$, their peak emission is also redshifted into the near-infrared which makes the both JWST Near Infrared Camera (NIRCam) and Roman Space Telescope Wide-Field Instrument (WFI) ideal for observing them. This makes it a possibility to serendipitously observe an ongoing TDE in deep field JWST imaging. In this work, we use the COSMOS-Web survey, the widest area deep JWST survey (Casey et al., 2022), to search for potential high-redshift TDEs. Detections of TDEs at $z>4$ could give valuable insights into SMBH formation and growth. They would indicate within which galaxies lower-mass SMBHs exist in the early universe, potentially distinguishing between SMBH seed models (Volonteri et al., 2021).
Previously, Kochanek (2016) has predicted that high-redshift TDEs would not be observable in ground-based wide-field time domain surveys. Here, we extend to high redshifts by leveraging deep space-based imaging without the use of time series. We assert in a companion work Karmen & et al. (in prep.) that, given the observed evolution of galaxy and SMBH properties with redshift, the TDE rate at high redshifts is likely greatly enhanced. As a result, in upcoming surveys it may be possible to study populations of TDEs at high redshift. Inayoshi et al. (2023) predicts the properties of high-redshift TDEs embedded within gas-rich AGN environments, in which the TDE occurrence rates are enhanced due to their compact and dense environments, compared to typical TDEs which occur in post-starburst galaxies. In this work, we complement this by a search exclusively for TDEs in galaxies not containing luminous AGN in order to utilize observed properties from low-redshift TDE searches, which exclude all AGN-hosting galaxies.
Throughout this work we adopt a Planck cosmology (Planck Collaboration et al., 2020). We report optical magnitudes in the AB system. We correct observed photometry for galactic extinction using the Cardelli et al. (1989) extinction law, assuming $R_{V}=3.1$ and the Schlafly & Finkbeiner (2011) extinction map.
In this work, we report the discovery of a $z=5.02^{+1.32}_{-1.11}$ TDE candidate in the COSMOS-Web survey (Casey et al., 2022). In Section 2, we present a new methodology for identifying TDE candidates in deep imaging using colors and morphology. We compare to potential contaminants, and pose methods of confirming candidate TDEs. In Section 3 we apply this methodology to the COSMOS-Web survey data, and use historic imaging of the COSMOS field to evaluate the potential variability of objects we select. We present the properties of HZTDE-1, a transient high-redshift TDE candidate in Section 4. We compare potential supernova models for this object to a TDE explanation. We discuss the need for spectroscopy of this object, and its implications if validated in Section 5. We summarize our work in Section 6.
[…]
3
2 Identifying High-Redshift TDEs in Deep Surveys
In this section, we model the signatures of TDEs similar to the local population, and develop a method for identifying them in wide field infrared surveys using the colors and morphologies. To first order, the UV/optical SED of a TDE is well-approximated by a black body with a typical temperature of $\sim 10^{4.3\pm 0.15}~{}K$ (van Velzen et al., 2020, 2021; Hammerstein et al., 2023; Yao et al., 2023). We simulate TDEs with the the same temperature distribution as the latest sample of spectroscopically confirmed TDEs from ZTF (Yao et al., 2023), and simulate their volumetric density using the ZTF $g$-band luminosity function. TDEs typically do not cool as they fade, so we assume constant-temperature blackbodies for their entire evolution. Their light curves are observationally well-parameterized in van Velzen et al. (2021) as follows:
$L_{\nu}(t)$ $=L_{\nu_{0}\,\rm peak}~{}\frac{B_{\nu}(T_{0})}{B_{\nu_{0}}(T_{0})}$
×$\times\begin{cases}e^{-(t-t_{\rm peak})^{2}/2\sigma^{2}}&t\leq t_{\rm peak}\\ e^{-(t-t_{\rm peak})/\tau}&t>t_{\rm peak}\\ \end{cases}$ (1)
with the luminosity in the $g$-band in erg s-1 $44.6842.68<\log(L_{g})<$ 44.68, the Gaussian rise time $1.30.4<\log(\sigma)<$ 1.3 days and the exponential decay time $2.31.2<\log(\tau)<$ 2.3 days from ZTF. We allow TDEs to fade until they reach the magnitude limit of the filter/survey observing them, to account for the time evolution of the TDE.
2.1 The Host model
We base our predictions of properties of TDE host galaxies on local observations of TDE hosts. It has been observed that TDE hosts are disproportionately in the green valley with compact nuclei (Hammerstein et al., 2021), and may show a preference for post-starburst galaxies (French et al., 2016). Because of this, we base our host galaxy models on the total sample of 30 ZTF TDE hosts from (Hammerstein et al., 2021) to capture their observed properties. To extend the observed host spectra to their rest-frame UV, we use Prospector to model their spectra and perform synthetic photometry (Johnson et al., 2021). We use the parameters from the original fits of the ZTF TDE hosts in Hammerstein et al. (2021), which are estimated by maximum a posteriori fits of Flexible Stellar Population Synthesis models (Conroy et al., 2009) to broadband photometry from SDSS, Pan-STARRS, and GALEX. When calculating properties of TDE hosts, we simulate every known TDE host to account for their full diversity of properties. When simulating observed TDEs in a survey, we include the host spectrum by adding a mean template, derived from the Prospector models, to the TDE spectrum before performing synthetic photometry.
We also use these ZTF TDE hosts as a basis for the physical sizes for typical TDE hosts. We query SDSS photometry for the TDE hosts in ZTF, and convert their angular sizes to physical sizes (Blanton et al., 2017; Hammerstein et al., 2023). We find that ZTF TDE host galaxies contain 90% of petrosian flux in a diameter of $8\leq D_{90}\leq 20$ kiloparsecs. This corresponds to range of angular diameters between $1.4$ and $4.7$ arcsec at $z=8$. This angular size will increase at even higher redshifts due to the decrease in angular diameter distance with increasing redshift beyond $z\sim 1.5$ (Hogg, 1999). As NIRCam has a resolution of 0.07 arcsec at 2 microns, all TDE host galaxies above the survey magnitude limit will be resolvable as an extended source in COSMOS-Web. At early epochs ($z\sim 7$), galaxies are expected to be intrinsically smaller and fainter (Grazian et al., 2012; Kawamata et al., 2018), such that even TDEs in unusual hosts will likely be point sources. Because of this, we predict that the majority of TDEs at high redshifts will outshine their hosts in the COSMOS-Web filters and be point sources.
Because all of the host galaxies we observe are at $z<1$, the galaxies we use are of higher mass and more luminous than those which we expect at high redshifts. High redshift galaxies are expected to be, and now observed to be, primarily lower-mass, blue, star-forming galaxies. While there is a minority of these galaxies which are exceptionally luminous without SMBHs above the Hills mass due to extremely active star formation (e.g. Franco et al., 2024; Casey et al., 2024; Castellano et al., 2024), it would be extremely difficult to detect a TDE in one because its star formation is bright and blue, making the TDE emission a sub-dominant component of its UV SED. Therefore, we do not consider them in our analysis. This makes the local TDE hosts we observe effectively upper limits for higher-redshift host galaxy luminosity. In summary, TDEs trace lower-mass SMBHs which trace, in turn, lower mass galaxies. These are expected to be even less massive at high-redshifts, meaning we expect the majority of TDEs to be hostless.
2.2 Dust extinction
Interstellar dust at high redshifts is something that is hotly debated and poorly understood. To first principles, high-redshift galaxies should have very low metallically, and should therefore have less dust attenuation than galaxies today (Fisher et al., 2014). This effect is observed in local metal-poor galaxies (Madden et al., 2013), and is beginning to be seen in JWST observations of high-redshift metal-poor galaxies as well (Heintz et al., 2023). Furthermore, preliminary observations of bright $z>10$ galaxies in the first JWST surveys can be explained by very low levels of dust (Finkelstein et al., 2023).
UV/optical TDEs are not typically observed in dusty galaxies (French et al., 2020; Hammerstein et al., 2021). This may be due to selection effects, or may be due to intrinsic host galaxy preferences. Our local volumetric TDE rate primarily includes TDEs which occur in local, quiescent galaxies which will decrease in number density with redshift. Combining this effect with high-redshift dust reduction, we expect that dust attenuation will have a small effect on the appearance of TDEs at high redshift. Despite this, we still include varying levels of dust in our modeling, simulating $A_{V}$ up to $1.2$, which is the dustiest any observed TDE host is in Hammerstein et al. (2023) fits to host photometry. Furthermore, the observation of a TDE implies the existence of a line-of-sight to the TDE emission which is not obscured by dust. Dust attenuation therefore affects the overall rates of observed TDEs more than the emission from the TDEs which are observed.
2.3 High Redshift TDEs
Figure 1: Example spectra of modeled $15000$ K TDEs at maximum and their host galaxies at four different redshifts: $z=1$ (dark blue), $z=2$ (light blue), $z=4$ (orange) and $z=6$ (dark red). Solid lines represent the TDE spectrum, while dashed lines show the host galaxy spectrum. The COSMOS-Web magnitude limits for each NIRCam filter are indicated in gray. At higher redshifts, host galaxies typically drop below detection thresholds, leaving TDE emission as the dominant observable component.
$C_{1}$ color $C_{2}$ color $a_{1}$ $a_{2}$ $a_{3}$ Min $C_{1}$ color Max $C_{1}$ color Min $C_{2}$ color Max $C_{2}$ color
(mag) (mag) (mag) (mag)
F115W - F [^4] W F150W - F277W $0.0024$ $0.4766$ $0.0473$ $-1.4$ 3.5 $-1.1$ $1.5$
F115W - F150W F277W - F [^4] W $0.0242$ $0.5215$ $-0.6974$ $-0.4$ $1.3$ $-1.0$ $0.1$
F115W - F150W F150W - F277W $0.0605$ $1.3533$ $-0.5377$ $-0.4$ $1.3$ $-1.1$ $1.5$
F150W - F277W F277W - F [^4] W $0.0005$ $0.3859$ $-0.4900$ $-1.1$ $1.5$ $-1.0$ $0.1$
Table 1: Parabolas of best fit to the high-redshift TDE color-color relationship, as seen in Figure 3 anr parameterized in Eqquation 2. Each line fit is a parabola to accurately model their slight curve in color-color space. These fits are no longer valid when there is significant Lyman dropout in the F115W band.
We find that, given our simulated TDEs and assumptions about their typical host galaxies, most high-redshift ($z>4$) TDEs will be point sources with unique colors. At $z>4$, a typical TDE host galaxy drops in apparent magnitude below the detection limits of the COSMOS-Web survey, as seen in Figure 1. A TDE near peak brightness is visible to beyond $z\sim 10$. The SED of a high-redshift TDE is dominated by its blackbody emission, which is redshifted from the rest-frame UV to the observer-frame near-infrared NIR. Because a redshifted blackbody is another blackbody, there is a degeneracy in color-color space between temperature and redshift, meaning all idealized TDEs lie upon a nearly straight line in 2D color-color space (slightly curved due to relativistic Doppler shift and dust extinction). In our selection for high-redshift TDEs, we restrict to a region that follows the color-color relationship in any given color combination. This is shown as the three dark purple lines outlining the TDE region in Figure 3. We also fit bounds to each individual color, which, together with the parabola fits, are presented in Table 1. We fit a parabola to the color-color relationship of TDEs in each filter combination, parameterized as:
$C_{2}=a_{1}C_{1}^{2}+a_{2}C_{1}+a_{3}$ (2)
where $C_{1}$ and $C_{2}$ are two colors. The values of coefficients $a_{1}$, $a_{2}$, and $a_{3}$ for each pair of colors are listed in Table 1, alongside the bounds on $C_{1}$ and $C_{2}$ form our simulations. Additionally, we fit the minimum apparent magnitude expected for a TDE at any redshift of a given color, shown in Figure 3, as
$m_{F115W}=2.15C_{F115W-F150W}+24.2$ (3)
To check our blackbody idealization, we take the following TDEs that have Hubble UV spectroscopy, and simulate them at high redshifts: iPTF16fnl, ASASSN-15oi, AT2022dsb, AT2018bsi, and ASASSN-14li (Charalampopoulos et al., 2022; Cenko et al., 2016; Blagorodnova et al., 2019; Brown et al., 2018; Hung et al., 2019; Hoogendam et al., 2024). We extend the UV spectrum into the optical ($>2600$) with a blackbody with the temperature fit by Mummery & van Velzen (2024), redshift the TDE spectra, and add in Lyman absorption shortward of rest-frame 1216Å (Inoue et al., 2014). We show these templates in Figure 2. These TDEs have colors centered on the relationship predicted, and scatter due to emission and absorption features, as well as continuum excess farther into the UV. This scatter, typically $\sim 0.2$ mag, defines the region near our idealized color-color relationships in which we expect actual TDEs to exist (the shaded region on the figure).
While we create these selection criteria for the COSMOS-Web survey, the same process can be done for any deep, wide-field infrared survey. Of particular interest is the Roman High Latitude Wide Area Survey (HLWAS) (Spergel et al., 2015). While the High Latitude Time Domain Survey (HLTDS) is designed to identify transients (Rose et al., 2021), the HLWAS will serendipitously observe hundreds of ongoing TDEs. Using similar methodology, one could select high-redshift TDEs in the HLWAS based on their colors and morphology, and pursue follow-up to confirm them. This search serves as a first test of such a method, for application in future wide field infrared surveys. For more intermediate-redshift hostless TDEs, this method may be possible to implement in optical wavelengths with the upcoming Vera Rubin Observatory co-adds and deep drilling fields (Ivezić et al., 2019).
Figure 2: Example templates used to test the TDE color-magnitude cuts, shifted to the rest frame. The red line is the HST spectrum, purple dotted is the blackbody used to extend the spectrum, and photometry is in blue. We switch from the HST spectrum to the blackbody at rest-frame wavelength $2600$, which is shown in the changing opacity of each line. We simulate this spectrum at higher redshifts to test our color cuts.
2.4 Alternative TDE Models
By extrapolating the blackbody NUV/optical spectrum of a TDE to the UV, we predict the spectrum of TDEs at rest-frame wavelengths at which their spectra are relatively poorly constrained, between $900$ and $1500$. Furthermore, the presence of Bowen lines in TDE spectra (e.g. Leloudas et al., 2019) implies that the flux predicted by a blackbody at $228$ is $\sim 5$ orders of magnitude less than what is needed in order to ionize He II and cause Bowen fluorescence. In the intermediate wavelength that we observe, between $900$ and $1500$, a blackbody may also under-predict the TDE flux.
To account for this potential discrepancy, we also consider a TDE spectrum with enough UV flux in order to trigger Bowen lines. For this test, we consider the Dai et al. (2018) model, which uses 3D GRRMHD simulations to simulate a super-Eddington accretion disk, and Monte Carlo radiative transfer to simulate reprocessing; we do not claim that this model is correct but treat it as an example of a spectrum which is extremely luminous in the UV. We select emission from a viewing angle between 22.5 $22.5-\>45^{\circ}$ from a simulated spectrum which matches a blackbody spectrum of a few ×$\times 10^{4}$ K in the optical. This model is not used for any rate calculations or final conclusions, but is a comparison point for TDE colors and identification methodology.
We find that the Dai et al. (2018) model falls well within our color-color and color-magnitude bounds, typically within $0.1$ mag of the central TDE color-color lines that we fit. While it would skew our fit to that of a hotter or lower-redshift TDE, spectroscopic follow-up would confirm our TDE redshift regardless of the underlying continuum model.
2.5 Contaminants
We model any point sources which may have similar colors to high-redshift TDEs to verify that high-redshift TDEs lie in a unique region of color-color and color-magnitude space. This varies dramatically by survey, but we demonstrate these color properties for COSMOS-Web, as the majority of future wide-field deep surveys will be in the near-infrared. We show examples of these color-color and color-magnitude relationships in Figure 3. The Roman HLTDS and HLWAS may in fact be easier to identify high-redshift sources in, because optical imaging from the surveys means that a Lyman break will be observed to verify the redshift of a source. This breaks multiple degeneracies we see in the JWST NIRCam infrared colors. The contaminants we model are as follows:
Lower redshift supernovae: SNe Ia, SNe IIP and SNe IIn at $z>2$ may be point sources in JWST data. We specifically model SNe which are point sources, so we do not consider contribution in their colors from host galaxy light. We predict that no SNe Ia will be detected beyond $z>5$ in COSMOS-Web. Already, supernovae with host redshifts from 0.6461 to 3.6 have been discovered using NIRCam imaging in the JWST Advanced Deep Extragalactic Survey (JADES) (Pierel et al., 2024; Siebert et al., 2024; DeCoursey et al., 2024; Coulter et al., 2025). To model these, we use from SNANA SNe II models (Kessler et al., 2009) of observed SNe at their peak luminosities, which have been implemented in sncosmo. Our SN Ia models are SALT3 models which extend into the NIR (Pierel et al., 2022), with parameter sweeps over $-2\leq x_{1}\leq 2$, $-0.3\leq c_{1}\leq 0.3$, and $z\leq 5$, simulated at phases $-20\leq\textrm{p}\leq 50$ days. The SNANA have only an amplitude parameters, so we sweep over a range of luminosities for each template. We set bounds on each supernova subtype’s luminosity distribution based on observed samples, $-16<M_{b}<-19$ for SNe Ia (Phillips, 1993), $-16\leq M_{V}\leq-19.5$ for SNe Ib/c (Drout et al., 2011), $-16\leq M_{b}\leq-18.5$ for SNe IIP (Sanders et al., 2015), $-13\leq M_{g}\leq-22$ for SNe IIn (Nyholm et al., 2020), and $-20\leq M_{g}\leq-22$ for SLSNe (Gomez et al., 2024).
We find that SNe Ia are sufficiently bluer than TDEs at lower redshifts that match their apparent magnitudes, and there is no confusion at any redshift. SNe Ib/c can indeed match the colors and magnitude of high-redshift TDEs at redshifts $z<1$, however their hosts are expected to be detected in COSMOS-Web and previous deep optical imaging of the COSMOS field. SNe II are even bluer, and can pass as $z>4$ TDEs around $z\sim 2$, but would need rest-frame $b$-band absolute magnitudes $\lesssim-20$, which is not observed in SNe IIP. SNe IIn can become this luminous at maximum brightness, serving as our main contaminant if their hosts are not detected at $z\sim 2$.
Superluminous supernovae (SLSNe): SLSNe, alongside TDEs and pair-instability supernovae, will be among the few transients which are visible up to very high redshifts. We model them using the templates constructed from the Dark Energy Survey (DES) SLSNe, which have a blackbody continuum with UV absorption features superimposed (Angus et al., 2019). Because their absorption strengths evolve and the SLSNe cool after their explosion, we simulate them up from $-10$ to $+35$ rest-frame days from peak with early-time temperatures between $5000$ K and $18000$ K. We use sncosmo to model these and perform synthetic photometry (Barbary et al., 2016). Again, we do not consider contribution from the host galaxy.
SLSNe are persistent contaminants given the low-mass nature of their typical hosts, and their colors and luminosities which significantly overlap those of TDEs (Schulze et al., 2018). A candidate TDE of a given high redshift could be a SLSN of comparable redshift. Time domain information, or spectroscopy is needed to distinguish a candidate TDE from a SLSN: SLSNe typically cool as they fade, whereas TDEs remain constant in color. A fading source with no color changes in two to three total observations over rest-frame months (observer-frame 66668888 months between observations) is an extremely strong TDE candidate. Finally, spectroscopy can identify typical TDE emission lines, such as broad H-alpha and He II, which can indicate a TDE (Gezari, 2021). However, a class of spectral “featureless” TDEs also exists, which would not have any strong emission line features (Hammerstein et al., 2023). Therefore, a combination of time-domain and spectroscopic monitoring is necessary for this rare case.
Quasars/AGN: We model both Type 1 and Type 2 AGN for comparison. We use the following SWIRE Type 2 AGN templates: QSO2, Torus and Sey2 (Polletta et al., 2007). We use the templates as they are integrated into the pysynphot Python package (STScI Development Team, 2013), and renormalize them to the $M_{UV}=-19.5$ mag. This is a typical absolute magnitude for a high-redshift obscured AGN given the JADES narrow-line AGN sample (Scholtz et al., 2023). We find, however, that our selection is relatively insensitive to the luminosity function of Type 2 AGN, and depends on their SED shape alone. Our Type 1 AGN model is a composite quasar template which combines the SDSS quasar spectral template in (Vanden Berk et al., 2001) with NIR quasar spectral template constructed from IRTF spectra in Glikman et al. (2006). We calibrate the absolute magnitude of the model in the $r$-band to the high-redshift ($z=7$) QSO luminosity function in Matsuoka et al. (2023), with a break magnitude of $-25.6$ mag. We redshift this quasar model at redshifts up to z=8, just beyond the highest-redshift confirmed quasar, J0313–1806 (Wang et al., 2021).
Type 2 AGN are universally too red to be contaminants in our TDE sample. A $z>8$ type 1 AGN can be similarly blue and have comparable apparent magnitude as a $z\sim 5$ TDE, but the power-law slope of its spectrum can be clearly distinguished from a blackbody given four photometric filters. Neither serves as a major contaminant in our sample.
Stars: We implement the PHOENIX stellar templates for stars with $5.50<\log{g}<$ 5.5, $240002000<T_{\textrm{eff}}<$ 24000 and $-2<\textrm{Fe}/\textrm{H}<1$ (Husser et al., 2013). We use pysynphot both to generate their spectra and to perform synthetic photometry in the NIRCam filters. We later use the same PHOENIX models to carefully simulate cool dwarf stars, such as M and L dwarfs. We simulate cool dwarfs to Milky Way halo distances, up to $20$ kpc, although distances $>5$ kpc are unlikely due to the scale height of the Milky Way, or until they are too faint for detection in COSMOS-Web.
Ultracool dwarf stars with temperatures $2000\leq T_{BB}\leq 3200$ K can masquerade as high redshift TDEs because they have predominantly thermal cool continuum emission which resembles redshifted hot thermal emission; e.g. a 2400 K local blackbody has the same spectral shape as a $z\sim 6$ 15000 K blackbody. While the dwarf star spectral templates we have are slightly flatter than blackbodies, they become a contaminant when searching for higher-redshift ($z>6$), cooler ($<20000$ K) TDEs. To distinguish these, we need to detect either time evolution, or a Lyman break to verify a candidate as an extragalactic transient. This is most prevalent when selecting bright ($>25$ mag) sources, but a halo star could hypothetically mimic a faint TDE candidate.
Brown dwarfs: We use the Sonora Bobcat models (Marley et al., 2021). These are both rare, and they do not match colors of TDEs so this comparison is just for verification of their difference.
Point-like galaxies: A small fraction of very-high-redshift galaxies observed in JWST are compact enough to render them point-like. We model these using the JAGUAR JWST Extragalactic Mock Catalog (Williams et al., 2018). This catalog models the evolving luminosity functions, spectroscopic properties, and morphologies of galaxies up to $z\sim 15$, and can reproduce observations up to $z\sim 10$. We specifically use simulations which use templates of expected $155<z<$ 15 galaxies. We select galaxies which will be point sources by selecting those which have an observed angular size equal to or less than the FWHM of the F115W PSF, $0.07$ arcsec. We use the pre-synthesized photometry from these catalogs for a wide range of galaxy types and redshifts, up to $z=15$.
We find that $>99\%$ of unresolved galaxies simulated will be too faint, given their color, to resemble a TDE as seen in Figure 3. However, rare unresolved extremely UV-bright galaxies are possible (e.g. Topping et al., 2024). In cases of excess UV emission, a flat optical continuum is still typically observed in these galaxies. A nondetection in F770W is a strong indicator that the candidate is likely not a galaxy. However, it is still valuable to obtain a spectrum or time-domain information to verify a TDE.
Figure 3: Color-color and color-magnitude plots of high redshift ($z>4$) TDEs compared to potential point source contaminants, modeled as described in Section 2.5. The purple central line is the parabola fit (Equation 2 and Table 1) to the TDE color-color relationship, and the region surrounding it is the $0.2$ mag of scatter seen in the TDE template colors from section 2.3. The unshaded regions contain 95% of contaminants of a given type. The red star is HZTDE-1, the high-redshift TDE candidate we identify in Section 3. The black x’s are the synthetic photometry of our TDE templates at $z=5$. HZTDE-1 is notably brighter than these templates because the majority of the template UV spectra are taken weeks to months after maximum brightness.
To summarize, our main persistent contaminants are very bright compact blue point source galaxies, local M- and T-dwarf stars, and hostless superluminous supernovae. All three contaminants can be ruled out spectroscopically. To rule out a galaxy contaminant photometrically one can use a Lyman break, and compare the absolute magnitude of the source to that of a TDE. To rule out a star contaminant photometrically, one only needs to observe the existence of a Lyman break. To rule out a superluminous supernova photometrically a second epoch of imaging is needed to observe the decline of the source.
[…]
3
3 COSMOS-Web Search Methodology
COSMOS-Web is an ideal first test for this methodology. The COSMOS-Web survey images $0.54$ deg2 of the sky with NIRCam to depths of $27.45$ mag in F115W, $27.66$ mag in F150W, $28.28$ mag in F277W and $28.17$ mag in F [^4] W, and occasionally in the Mid-Infrared Instrument (MIRI) band F770W to $\sim 26$ mag (Casey et al., 2022). Decades of deep multiwavelength observations of the COSMOS field means that, for a given source, one can look back over 20 years of observations (e.g. Scoville et al., 2007; Capak et al., 2007; Koekemoer et al., 2011). Furthermore, the discovery of a high redshift TDE in COSMOS-Web could be evidence of the high-redshift TDE rate enhancements that we outline in Karmen & et al. (in prep.). Given a constant TDE rate in the entire universe, Karmen & et al. (in prep.) predicts $0.7$ TDEs in the COSMOS-Web dataset. However, given early-universe TDE rate enhancements discussed in Section 5.1 and in Inayoshi et al. (2023), this rate could be much higher. We apply our method to COSMOS-Web catalog (M. Shuntov, L. Paquereau et al. in prep.; see Akins et al. (2024) for details on catalog construction): we simulate TDEs in the NIRCam filters F115W, F150W, F277W and F [^4] W using the methods described in Section 2, apply cuts based on their colors, magnitudes, and morphologies, and then check in previous observations to verify if these sources are indeed transient.
To select point sources, we calculate the ratio between the flux in a $0.5$ arcsecond aperture and in a $0.25$ arcsecond aperture. We select stars by LePHARE template fitting (Arnouts & Ilbert, 2011; Ilbert et al., 2006) to their SED shapes and apparent magnitudes, and find the aperture flux ratios for these sources. We calibrate our point source flux ratios to the range of flux ratios for stars, as shown in Figure 4. We used the bounds of this relationship, the upper and lower limits on the flux ratio for a given magnitude, to select point sources in COSMOS-Web. We restrict only to objects which are point-like in F115W and F150W, the filters in which we most confidently expect no galaxy light beyond $z\sim 4$.
Figure 4: Flux ratio between a $0.25"$ aperture and $0.5"$ aperture in F150W vs magnitude for all sources in the COSMOS-Web catalog (gray points). Sources which are classified as stars via LePHARE SED fits (blue points) are used as a point source selection guide. The black dashed lines show the region within which we select point sources.
Once we have selected point sources, we use our simulated TDEs (Section 2) to define the region of color-color and color-magnitude space that high-redshift TDEs occupy. We cuts in color-color and color-magnitude relationships to restrict only to the regions where we believe high-redshift TDEs occupy. This is shown in the purple region in Figure 3.
3.1 Cross-match with previous COSMOS data
The COSMOS field has been repeatedly surveyed for the past 20 years, since the original Hubble Space Telescope COSMOS survey, which were carried out between October 2003 and November 2005 (Koekemoer et al., 2007). We use this original catalog to check the transient nature of our candidates. The HST COSMOS survey was conducted using the F814W filter, which does not overlap in wavelength with our NIRCam observations but allows an initial check for variability.
The COSMOS field also has deep optical imaging with the Hyper Suprime Cam (HSC) on the Subaru telescope, originally in Capak et al. (2007), and later in (Tanaka et al., 2017) to $5\sigma$ AB magnitude depths of $27.8$ in the $g$-band, $27.7$ in the $r$-band, $27.6$ in the $i$-band, $26.8$ in the $z$-band, and $26.2$ in the $y$-band. These observations were conducted between March 2014 and May 2015, giving them an approximately decade-long observer-frame cadence between both the COSMOS HST ACS F814W observations, and the COSMOS-Web survey. These later HSC data comprise the COSMOS2015 catalog (Laigle et al., 2016)
In addition, there is previous near-infrared imaging taken in $YJHK_{s}$ to $\sim 26$ mag and in the narrow band $NB118$ to $24.6$ mag taken from the UltraVISTA survey, with the VIRCAM instrument on the VISTA telescope (McCracken et al., 2012). We use stacks of UltraVISTA data observed between December 2009 and June 2016. There are previous mid-infrared data from Spitzer/IRAC channel 1, 2, 3, 4 images from the Cosmic Dawn Survey (Euclid Collaboration et al., 2022), which are stacks from all data taken up to January 2020. However, the IRAC data does not have sufficient angular resolution to distinguish very small, crowded sources so we cannot often use it.
The COSMOS2020 catalog (Weaver et al., 2022) contains all of the data mentioned (with a summary of photometry limits in Weaver et al. (2022) Table 1), and is publicly archived. In this analysis, we use forced photometry for previous surveys at the location of a detection in COSMOS-Web, and classify a nondetection as any source with a $<3\sigma$ detection in flux. This is useful when a TDE candidate is sufficiently bright such that a non-detection in shallower ground-based data is constraining; the detection image in COSMOS2020 is a $izYJHK_{s}$ stack. We use these data to extend SEDs when visualizing candidates, but do not use them in our candidate selection or fitting in case of variability between the HSC observations and COSMOS-Web.
We search for our candidates from COSMOS-Web in the COSMOS2020 catalog using the IPAC COSMOS Archive (COSMOS Project, 2020) for detections in existing COSMOS catalogs within a matching radius of 1”. For the subset of candidates which are not in COSMOS2020, we download their cutouts and manually inspect them. We separate these cutouts into two groups: objects that have NIRCam F115W photometry that indicate they are above the HST F814W COSMOS magnitude limit of $27.8$ mag, and those expected to be below the magnitude limit. In particular, we are interested in objects that would be bright enough to be detected in the COSMOS2020 but did not find. If a source is detected in the HST 814W mosaics, we assume that is is not a transient source. However, we do not consider a nondetection in HST confirmation of its transient nature. Although we are searching for high-redshift transients that are significantly time dilates, a $z=8$ TDE would be bright for about 3 observer-frame years, so a $\sim 20$ year baseline should be enough for any high-redshift TDE to rise or fade.
3.2 Search Results
Figure 5: Left: Color-color diagram of all sources in black, all point sources in gray, and all selected candidates in blue points. TDE candidate HZTDE-1 is overplotted as the red star. Right: Color-magnitude diagram of all point sources in gray, and all selected candidates in blue points. TDE candidate HZTDE-1 is again overplotted as a red star.
We ultimately selected 1764 sources that passed our point source cuts, 117 passed our color-color, and magnitude cuts. We visualize these candidates in color-color and color-magnitude space in Figure 5. Among these sources, 82 of them had no listing in the COSMOS2020 catalog. We divide these into “bright” and “faint” sources. “Bright” sources are sources with $M_{F115W}>27.8$ AB mag, where $27.8$ mag is the detection limit in HST F814W in COSMOS2020, the deepest imaging of this field. These filters do not overlap– the $\lambda_{\rm max}$ of F814W is $9643$ and the $\lambda_{\rm min}$ of F115W is $9975$– but they are near enough to where nondetection in F115W approximates a nondetection in F814W. Therefore, “bright” sources should have mostly been detected in HST. We find that our candidates contain 79 “bright” sources and 3 “faint sources”. For all 79 sources without HST catalog entries, we download the HST F814W cutouts, and search through them by hand for nondetections. We ultimately determine by visual inspection that 72 of the 79 “bright” sources which pass our cuts do have a counterpart in F814W images, but are excluded from COSMOS2020. We note that these sources often lie on the edge of a tile, in the masked region near a bright star, or have inaccurate astrometry. Among the sources not seen in the manual inspection, six are very red and not bright in F184W. We are left with one source which is detected in COSMOS-Web but not in COSMOS2020. The remaining candidate peaks at $\sim 25$ mag in F150W, which is $>1$ mag above the limit of previous infrared imaging with UltraVISTA. This source is not detected in UVISTA data to even $1\sigma$, confirming its nature as a transient and viable high-redshift TDE candidate. We will refer to this TDE candidate at HZTDE-1.
Figure 6: Top: Example of a high-redshift TDE candidate which we believe to be a $z=3$ galaxy contaminant. Despite having a blackbody-like SED shape in NIRCam, the HSC photometry shows that the SED is flat in the optical, with a Lyman break beginning in the g-band. Its F770W fluc greatly exceeds that of a blackbody. The source appears point-like in F115W and F150W, but in the F277W imaging it can be seen to be slightly extended, although compact. The best-fit galaxy template is plotted in red, star template in light blue, and quasar in purple. Bottom: Example of a high-redshift TDE candidate which we believe to be a local cool dwarf star. The source has no listing in the Hubble Source catalog, although it is detected in COSMOS2020. This source has no Lyman break, and is detected to $5\sigma$ in every HSC band.
These candidates are selected only by JWST photometry, and the HSC photometry is consistent with a blackbody with the same peak as a high redshift TDE for the vast majority of sources, as seen in Figures 6 and 7. Figure 7 shows an example fit of our simplified high-redshift TDE model to HZTDE-1, characterized only by absolute magnitude, blackbody temperature, and redshift, to the photometry. We find that, while the model fits well to the majority of our data, the degeneracy between redshift and temperature means that the fits are extremely dependent on the starting guess and there is no way to discern a source’s redshift without observing a Lyman dropout. Additionally, shown as a comparison, this could also match the SED of an M-dwarf star. Therefore, for these data it is essential that we utilize all previous HST imaging to check for variability and optical HSC data to look for a dropout.
Cross-identification in the Hubble Source Catalog suggests that candidates that have Lyman dropouts are likely high-redshift galaxies. In the HST COSMOS catalog, many of our faint candidates are labeled “high-redshift galaxy”, some with spectra confirming these labels. This is explained by our point source selection: we select point sources based on their flux ratios in two small apertures (See Section 3), rather than doing actual morphological fitting. This means that a sample of slightly lower-redshift, brighter and bluer galaxies could contaminate the sample by having an approximate angular size of order of the smaller aperture we use, rather than a smaller angular size than the NIRCam PSF. Better morphology selection could reduce this contamination, but nothing can eliminate rare perfect contaminant compact, blue, high-redshift galaxies other than time-evolution. We show an example of these high-redshift galaxy contaminants in Figure 6. HSC photometry confirms that it is indeed not a blackbody, and has a Lyman dropout in the g-band. Additionally, while it appears point-like in F115W and F150W, the F277W imaging shows that it is an extended source. The object is just under a $5\sigma$ detection in the HST image, but s visible in the cutout. We also search all of our objects in the Chandra X-ray imaging of the COSMOS field, and find no X-ray sources associated with any object (Civano et al., 2016). This is expected for a high-redshift TDE, because their soft X-ray spectrum would be completely absorbed by the CGM beyond $z\sim 3$.
Objects without Lyman dropout are likely cool dwarf stars. Stars with temperatures $2400\textrm{K}<T<3500\textrm{K}$ have blackbody peaks around 1.2 $\mu$m, the same observed wavelength as a $z>5$ TDE. These stars are abundant in the Milky Way, although are less common when looking out of the Galactic Plane. Figure 6 shows a source which we believe is a star. It has no listing in the COSMOS2020 catalog, but manual inspection can identify the corresponding object in previous COSMOS imaging.
4 Properties of the TDE Candidate HZTDE-1
Figure 7: Cutouts (top panels) and SED (bottom panel) of high redshift TDE candidate, HZTDE-1. The cutouts include imaging from UltraVISTA, Hyper Suprime Cam, Hubble ACS, and JWST NIRCam. The source is only detected to $>3\sigma$ in the NIRCam images. The SED shows the observed SED as red circles, synthetic photometry for the best-fit galaxy spectrum (purple line) and TDE spectrum (red line) in diamonds, and all photometric upper limits as downward-pointing triangles. Errorbars in magnitude exist for the NIRCam data but are too small to be seen. The transmission curve for each filter is plotted below it. A simulated spectrum of TDE 2019qiz at $z=5$ is overplotted as a dotted line, and a template for a $24,000$ K superluminous supernova at $z=3.6$ is the yellow line.
The TDE candidate HZTDE-1 is shown in Figure 7. This source was the only object which is a point source in NIRCam, has no optical counterpart, has the SED shape of a high-redshift blackbody, and does not appear in the COSMOS2020 catalog. Furthermore, HZTDE-1 is not detected in previous deep UltraVISTA imaging, which verifies the transient nature of the object. Due to possible contamination from a nearby galaxy, we verify the nondetection in HSC and UltraVISTA by performing aperture photometry to manually compare the flux at the location of HZTDE-1 to the background flux. We describe this procedure in Appendix A, and find that there truly is no detection above background in any previous optical or NIR imaging. LePHARE spectral template fitting does not settle on a good classification for this source given its SED: its reduced chi-squared statistic (per degree of freedom) for the best star template is $\chi_{\nu}^{2}=477$, for an AGN template is $\chi_{\nu}^{2}=1050$, and $\chi_{\nu}^{2}=401$ for the best galaxy template, which fits a galaxy at $z=7.33$. We also show in Figure 7 a composite spectrum of TDE 2019qiz, which is constructed by connecting an ultraviolet spectrum taken with HST UVIS (Hung et al., 2021) and an optical spectrum taken with the X-Shooter instrument on the Very Large Telescope (Nicholl et al., 2020). We select this TDE due to its broad wavelength coverage near maximum brightness; it best fits the HZTDE-1 photometry at $z=4.1$ due to its lower blackbody temperature ($15,000$ K), but we simulate its spectrum at $z=5$ to show the locations of common TDE emission lines.
We approximate its redshift through maximum a posteriori fitting of our blackbody TDE model with Gaussian uncertainties optimizing through Monte Carlo Markov Chain sampling. We utilize the ZTF TDE sample’s distribution of TDE temperatures and luminosities as priors. Our prior on redshift is a uniform distribution between $z=3.5$ and $z=7.5$. We estimate a lower bound on the redshift through our simulations of TDE hosts: we expect that the lowest-mass TDE host would not be detected in the two short-wavelength NIRCam filters of COSMOS-Web at $z\geq 3.5$. We can set a strict upper limit on the redshift because at $z=7.5$, the F115W emission would have been Lyman absorbed. We find our most likely TDE to have $z=5.02^{+1.33}_{-1.11}$, $M_{g}=-21.15^{+0.21}_{-0.13}$, and $\log(T_{BB})=4.31^{+0.09}_{-0.09}$. We show the MCMC sampling of the posterior distribution in Figure 8. If HZTDE-1 were as cool as the lowest-temperature known TDEs, $\sim 10^{4}$ K, it could be as low-redshift as $z=2$; however, as $z=2$ a typical $10^{10}~{}M_{\odot}$ host galaxy would be resolved in COSMOS imaging.
Figure 8: MCMC samples of the posterior distribution for each parameter in the TDE model fit to the HZTDE-1 photometry. Errorbars are wide due to degeneracies between temperature, redshift, and luminosity. We use the observed temperatures and maximum luminosities of ZTF TDEs as priors. We also set bounds on the redshift fit as decribed in Section 4.
4.1 Comparison to supernova models
There are three nearby galaxies that HZTDE-1 could belong to if it is a lower-redshift supernova. We assess its possible association with the galaxy directly above it in the cutouts shown in Figure 7, because it is the closest galaxy, with an angular separation of $0.998$”, the brightest galaxy with $M_{F444W}=23.7$ AB mag, and the most extended with effective radius $R_{e}=0.20$” as modeled by GALFIT (Peng et al., 2002). The GALFIT model for the sources in the field is shown in Figure 9, alongside the F [^4] W image and the residual. The best-fit galaxy spectrum to this source’s SED has a photometric redshift $z_{\rm phot}=1.75^{+0.1}_{-0.07}$. The SED is well-sampled from the HSC $g$-band to NIRCam F [^4] W, and the model spectrum fits accurately with a $\chi^{2}=1.72$ per degree of freedom. We follow the technique developed in Gupta et al. (2016) to assess host galaxy association for a given supernova through calculating the directional light radius (DLR), which is the elliptical radius of a galaxy in the direction of the supernova. They then calculate the $d_{\rm DLR}$, which is the distance between the supernova and the galaxy normalized by the DLR. They utilize both $HST$ ACS data and MICECATv2.0 simulations to show that any SN with a galaxy $>5$ DLR away should be assigned “hostless”, and achieve $\sim 95\%$ completeness with a similar cut. We calculate the unitless dDLR for HZTDE-1 to be $5.05$ relative to the nearest galaxy, making it highly likely to be a hostless transient. This dDLR is even higher for the short-wavelength NIRCam filters ($>6$ DLR in F115W), which have higher spatial resolution.
Figure 9: GALFIT modeling of the morphology of the sources in the F [^4] W imaging of HZTDE-1. We display the size of $1$ directional light radius (DLR) from the galaxy to the transient to show the accuracy of the modeling. The HZTDE-1 is $>5$ DLR from the galaxy, unlikely to be associated. In shorter-wavelength filters, this DLR separation increases.
We assume that HZTDE-1 is unlikely to be a “hostless” non-superluminous supernova. However, given the rare case that HZTDE-1 is associated with the nearest galaxy despite extreme distance, we fit supernova SEDs for each supernova subtype at its photometric redshift, $z=1.75\pm 0.10$. We use the full range of supernova properties described in Section 2.5. For every supernova subtype, we fit every available model in sncosmo (Barbary et al., 2016), extending to the infrared with a blackbody when the model has insufficient wavelength range. We do not fit models for Population III SNe, Pair-Instability Supernovae (PISNe), or other exotic transients. We show our best-fit supernova spectra in Figure 10, and their respective lightcurves in Figure 11. We find that HZTDE-1 is significantly too red to be a SN Ia at $z\sim 1.75$ and too bright to be a SN Ia at $z>2$. HZTDE-1 is $\gtrapprox 1$ too bright to be explained by a Ib/c at $z\sim 1.75$. The SED is somewhat plausibly explained by an anomalously bright IIP before maximum brightness or a bright IIn prior to maximum brightness, but both under-predict the flux in F [^4] W at $z\sim 1.75$, which would be $1.6~{}\mu$m in the rest-frame. These best models are shown in Figure 10. This excess flux could be explained by contribution from thermal dust emission, but given the extreme offset from the potential host it is unlikely that the supernova would be embedded in dust. In summary, for HZTDE-1 to be a supernova associated with its nearest galaxy it must be an anomalously bright SN IIn at an extreme offset with an infrared flux excess.
Figure 10: Modeled spectra and synthetic photometry of best-fit Type Ia, Type Ib/c, IIP, and IIn supernovae compared to the SED of HZTDE-1. All SNe are within $z=1.75\pm 0.1$, in order to be associated with the most likely host galaxy. Type Ia and Ib supernova cannot successfully fit the SED. A very bright IIP or IIn could fit the majority of the SED, but struggle to produce the same infrared luminosity.
There is a possibility that HZTDE-1 is a superluminous supernova (SLSN), as shown in Figure 7. SLSNe have similar intrinsic luminosities to TDEs and their spectral continua are primarily blackbody thermal emission (Gomez et al., 2024). Their host galaxies tend to be less massive, meaning that at large enough distances ($z\gtrsim 3$) they may also outshine their hosts and be point sources. Therefore, to be certain whether HZTDE-1 is a high-redshift TDE or high-redshift hostless supernova, either follow-up spectroscopy or monitoring of its photometric evolution is necessary. With a single photometric detection, the relative rates of TDEs and SLSNe at $z>3$ is the main factor that determines what HZTDE-1 likely is. Current best estimates of the volumetric rate of SLSNe place them somewhere between $0.91\times 10^{-7}~{}\rm{SNe}/\rm{year}/\rm{Mpc}^{3}$ at $z=1$ (Prajs et al., 2017) and $1.99\times 10^{-7}~{}\rm{SNe}/\rm{year}/\rm{Mpc}^{3}$ at $z=0.2$ (Quimby et al., 2013), as compared to the current best estimate for the local optical TDE volumetric rate of $3.1\times 10^{-7}~{}\rm{TDEs}/\rm{year}/\rm{Mpc}^{3}$. Tanaka et al. (2013) predicts the volumetric rate at SLSNe at higher redshifts (up to $z\geq 10$) by scaling it with the cosmic star formation rate (SFR) density using two SFR models (Robertson & Ellis, 2012; Hopkins & Beacom, 2006). They calibrate to the SLSN low-redshift rate, and account for the initial mass function and the fraction of massive stars to find that the SLSN volumetric rate increases with redshift until $z\sim 5$, where the SFR density peaks, and then the rate decreases with redshift indefinitely. At its maximum, $z\sim 5$, the volumetric SLSN rate is 10×10$\times$10 × the local rate. We expect the TDE rate to also be elevated by at least 10×10$\times$10 × the local rate, as shown in Karmen & et al. (in prep.). HZTDE-1 would need to be extremely blue for a SLSN, with a best-fit temperature $24,000$ K at best-fit redshift $z=3.6$, but this is possible briefly before peak and during peak brightness. While this makes a SLSN somewhat unlikely, given the comparable rates there is no way to rule out a SLSN without further photometric monitoring or spectroscopy.
Figure 11: Lightcurves of best-fit supernovae models to HZTDE-1, given that they are at redshift within $0.1$ of the closest galaxy. Each panel is in a single COSMOS-Web NIRCam filter. The best-fitting models, IIP and IIn, do not adequately explain the F [^4] W flux, which would be at $\sim 1.6\mu$m in the rest frame.
5 Discussion and Future Confirmation
In this work, we model the properties of TDEs at high redshift assuming that their populations are similar to those at low redshift. We argue that the assumptions made about their host galaxies (Section 2.1) and their colors (Section 2.3) are treated as bounds rather than constraints and that evolution of these properties with redshift will lie within our selection criteria. However, any number of these assumptions can be relaxed. In future work, for example, modeling of TDEs in extended hosts, or modeling of dust-extincted TDE emission can increase the efficiency of searches for TDEs in wide-field surveys. As this search offers an opportunity to study the UV emission shortward of $2000$ which is usually not observed, deviations from blackbody assumptions about TDE spectra, such as those mentioned in Section 2.4 may emerge from studies of high-redshift TDEs.
Both spectroscopic and photometric follow-up is needed to confirm HZTDE-1’s classification as a TDE and its redshift. TDEs have blue spectra with blackbody continua and broad Hydrogen and/or Helium emission features that narrow as the TDE evolves. A key distinction between TDEs and supernovae is their decline: TDEs fade slowly, typically as $t^{-5/3}$, and do not change temperature as they cool (van Velzen et al., 2020). Without imaging of a host galaxy to verify that HZTDE-1 is nuclear, combining identification these spectroscopic features and photometric monitoring is needed to confidently confirm it as a TDE. Spectroscopy, specifically in the infrared with JWST’s NIRSpec, is key to observing key TDE emission lines. X-ray emission is not expected because, as mentioned, the CGM will absorb all of the soft X-ray emission characteristic of TDEs at $z>3$. Longer wavelengths may be worth monitoring for later-time infrared flares due to dust echoes (van Velzen et al., 2016).
5.1 TDE rates in COSMOS-Web
In Karmen & et al. (in prep.), we calculate observed TDE rates as a function of redshift in LSST, Roman, and the COSMOS-Web survey. We find that if the TDE rate does not evolve as a function of redshift, there is a $68\%$ chance of observing a single TDE in COSMOS-Web, and a $7\%$ chance of observing a TDE at $z>4$. We show, however, that the observed properties of high-redshift galaxies would imply a dramatically enhanced rate of TDEs at high-redshift $105<z<$ 10. Namely, many galaxies are observed to be centrally compact (Finkelstein et al., 2023; Kartaltepe et al., 2023; Baggen et al., 2023; Guia et al., 2024), and may host nuclear stellar clusters (Ricotti et al., 2016; Sun et al., 2023) at high redshifts. This increases central stellar density, which in turn will lead to a ×$\times 100$ TDE rate enhancement (Kritos et al., 2024; Pfister et al., 2020). Additionally, galaxy mergers are shown to enhance TDE rates both observationally and theoretically. Host studies of optical/UV TDEs demonstrate a clear preference for galaxies which are believed to be products of mergers (French et al., 2020; Wevers & French, 2024), and the TDE rate is seen in hydrodynamical simulations to be enhanced by an order of magnitude for 100 Myr after a merger (Pfister et al., 2021). With a galaxy merger rate that increases as a function of redshift, peaking at cosmic noon (Ventou et al., 2017; Duncan et al., 2019; Perna et al., 2025) we predict TDE rates to be enhanced from mergers as well. In the past it was expected that TDE rates decrease with the SMBH mass function as a function of redshift. However, given the abundance of supermassive and overmassive black holes detected at early cosmic times (e.g. Maiolino et al., 2023; Larson et al., 2023; Kokorev et al., 2023), this phenomenon will be significantly reduced. TDEs are among the few luminous phenomena very close to an SMBH, and are believed to produce super-Eddington accretion at early times (Strubbe & Quataert, 2009) so their study at high redshifts can give an indication to the mechanisms by which early SMBHs accrete and grow.
The TDE rate can still be enhanced at high redshifts without the detection of an overabundance of TDEs in JWST imaging. Samples of infrared-bright TDEs find that the majority of TDEs with no detected optical/UV emission occur in star-forming galaxies (Masterson et al., 2024). It is known that the fraction of star-forming galaxies increases with redshift beyond $z>2$ (Madau et al., 1996). Therefore, if the majority of high-redshift TDEs occur in star-forming galaxies, it follows that their UV emission may be dust obscured, reprocessed, and emitted as rest-frame infrared emission. This possibility is supported in Inayoshi et al. (2023) which considers the enhancement of the rate of TDEs in obscured AGN. In order to detect the remainder of high redshift TDEs, if they reside in star-forming galaxies, one may need to look for the rest-frame infrared dust echoes. This rest-frame ultraviolet search is a first probe of these rate enhancements. In order to confirm the classification of HZTDE-1, spectroscopic follow-up must be done before the source completely fades.
6 Conclusion
In this work we search the COSMOS field for ongoing high-redshift TDEs, and successfully identify one $z\sim 5$ TDE candidate. We summarize our findings below:
  • We show that beyond $z>4$, given both the nature of TDE preferred host galaxies and the low masses of galaxies at high redshift, the high-redshift counterparts to local, optical/UV TDEs will be hostless point sources. We demonstrate that they will be k-corrected into the wavelength range of JWST’s NIRCam and could be detected to $z\sim 10$.
  • We develop criteria for identifying candidate $z>4$ TDEs given their NIRCam colors, comparing to contaminant stars, galaxies, AGN, and supernovae. For regions of potential contamination, we pose methods for distinguishing TDEs from these contaminants. This methodology can be applied to any deep field survey, and will yield a wealth of high-redshift TDEs from the Roman High Latitude Wide Area Survey.
  • We apply this selection method to the COSMOS-Web catalog, and identify 117 sources among $>700,000$ that pass our cuts. We systematically check these sources for lack of variability using the COSMOS2020 catalog, which compiles all previous imaging of the COSMOS field, and find one source which is a confirmed transient. We call this TDE candidate HZTDE-1.
  • We model HZTDE-1 as a possible supernova, both through SED modeling and host galaxy association. We find that HZTDE-1 is $>5$ directional light radii from the nearest galaxy, making an association extremely unlikely. We find that while the SED cannot be explained by a Type I supernova, a Type II (particularly IIn) at $z\sim 2$ can explain the SED mildly well, except for the farthest infrared emission. However, it is still possible that this transient is a “hostless” luminous or superluminous Type IIn supernova.
  • We model HZTDE-1 as a possible TDE, and find that the most likely TDE has an effective blackbody temperature of $\log(T_{\rm{BB}})=4.31\pm 0.09$, absolute g-band magnitude of $M_{g}=-21.15^{+0.21}_{-0.13}$ and redshift of $z=5.02^{+1.33}_{-1.11}$.
Follow-up observations of HZTDE-1, both for photometric fading and for spectroscopic information, are crucial for understanding its nature and confirming its redshift. If HZTDE-1 is a TDE, it could confirm an enhancement of the volumetric rate of TDEs in the high-redshift universe. As TDEs offer a probe of central SMBH mass independently from the virial methods that use AGN emission, a population of TDEs at high redshifts could offer insight into the nature of the SMBH mass function as a function of cosmic time. This, in turn, could place additional constraints on the methods by which SMBHs seed and grow.
7 Acknowledgements
The authors thank Christina Lindberg for discussions about dust, and Sean Carroll for his discussion of source selection. We also thank Muryel Guolo for his TDE wisdom and cynicism. This paper is based in part on observations with the NASA/ESA Hubble Space Telescope and James Webb Space Telescope obtained from the Mikulski Archive for Space Telescopes at STScI. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE2139757. The French contingent of the COSMOS team is partly supported by the Centre National d’Etudes Spatiales (CNES). OI acknowledges the funding of the French Agence Nationale de la Recherche for the project iMAGE (grant ANR-22-CE31-0007). This work was made possible by utilising the CANDIDE cluster at the Institut d’Astrophysique de Paris. The cluster was funded through grants from the PNCG, CNES, DIM-ACAV, the Euclid Consortium, and the Danish National Research Foundation Cosmic Dawn Center (DNRF140). It is maintained by Stephane Rouberol.
Appendix A Verification of optical nondetection
Because of the small angular distance between HZTDE-1 and the nearby galaxy, we wanted to verify that HZTDE-1 was not obscured by the galaxy in the lower-resolution UltraVISTA imaging. We use aperture photometry to test this as follows. First, in pixel space, we calculate the location of the centroid of the galaxy in each filter by fitting a 2D gaussian to the data in the photutils Python package (Bradley et al., 2025). Filter by filter, we find the angular separation between the galaxy centroid and the PSF-modeled coordinates of HZTDE-1 which we call $d$. Starting with the location of HZTDE-1, we place a series of apertures, each with a radius of $2$ pixels, at an angular distance $d$ from the galaxy, each offset from each other by a rotation angle of $20^{\circ}$, as shown in Figure 12. We calculate both the mean flux and the total flux within each aperture in UVISTA and find no increase in the flux at the location of HZTDE-1 compared to background in any filter. We then perform the same procedure using JWST imaging, with apertures of radius 3 pixels, as seen in Figure 12. As expected, we find a ×$\times 1000$ increase in flux relative to background in NIRCam, as seen in Figure 14.
Figure 12: Left: UltraVISTA $K_{s}$-band imaging of galaxy near HZTDE-1, surrounded by the apertures described in Appendix A. The red aperture is centered on HZTDE-1. Right: JWST NIRCam F150W imaging of the same galaxy and HZTDE-1, surrounded by apertures in the same location. Again, visibly, the red aperture contains HZTDE-1. For visualization purposes and because JWST has a very high angular resolution, the JWST apertures are slightly smaller.
Figure 13: Flux from UltraVISTA imaging in the apertures from Figure 12, with $0^{\circ}$ corresponding to the location of HZTDE-1. The left figure is the mean flux within each aperture for each filter, and the right figure is the total flux within each aperture. Neither has an increase in flux at the location of HZTDE-1 in any filter.
Figure 14: Aperture from JWST imaging using apertures in the same locations as for UltraVISTA relative to the nearby galaxy. For visualization purposes and because JWST has a very high angular resolution, the apertures are slightly smaller. $0^{\circ}$ corresponds to the location of HZTDE-1. Using NIRCam, there is a clear detection of HZTDE-1 with respect to the background.

### v4

@@
 In this work, we present the discovery of a high-redshift TDE candidate, HZTDE-1, in the COSMOS-Web survey with JWST’s NIRCam, using a novel selection technique based on color and morphology.
-We first outline a methodology for identifying high-z TDEs in deep infrared imaging surveys, leveraging their unique spectral energy distributions (SEDs) and morphologies of these transients.
+We outline a methodology for identifying high-z TDEs in deep infrared imaging surveys, leveraging their unique spectral energy distributions (SEDs) and morphologies of these transients.
+While focused on TDEs, this methodology could also be applied to find other UV-bright transients, such as superluminous supernovae (SLSNe).
 We apply this technique to COSMOS-Web in filters F115W, F150W, F277W, and F [^4] W, and identify HZTDE-1, a transient point source relative to archival UltraVISTA infrared observations.
-If we assume it is a TDE, we estimate its photometric redshift to be $z=5.02^{+1.32}_{-1.11}$.
+If we assume it is a TDE, we estimate a photometric redshift of $z=5.02^{+1.32}_{-1.11}$.
 HZTDE-1 cannot be explained by reasonable supernova or AGN models.
-However, we cannot rule out a superluminous supernova at $z\gtrsim 3$.
-If confirmed with follow-up observations, HZTDE-1 would represent the highest-redshift TDE discovery to date, and would suggest an enhancement of the TDE rate in the high-redshift universe.
+However, a SLSN at $z\gtrsim 3$ can also plausibly explain this transient and would be the highest-redshift known SLSN.
+If confirmed with follow-up observations, HZTDE-1 would represent the highest-redshift TDE discovery to date, and suggest an enhancement of the TDE rate in the high-redshift universe.
 Our method, which can be applied to future deep surveys with JWST and Roman, offers a pathway to identify TDEs at $z>4$ and probe black hole demographics at early cosmic times.
@@
 Unlike AGN, which prefer very massive SMBH hosts (Kauffmann et al., 2003), TDEs of main-sequence stars by non-rotating (Schwarzschild) black holes can only occur for lower mass SMBHs because the disrupting black hole has an upper mass limit of $M_{BH}<10^{8}M_{\odot}$ (Hills, 1975).
-Beyond this mass, the tidal radius $R_{t}=R_{\star}(M_{BH}/M_{\star})$ is within the Schwarzschild radius $R_{s}=2GM/c^{2}$.
+Beyond this mass, the tidal radius $R_{t}=R_{\star}(M_{BH}/M_{\star})^{1/3}$ is within the Schwarzschild radius $R_{s}=2GM/c^{2}$.
 The TDE rate’s sensitivity to the black hole mass make TDEs an excellent probe of the SMBH mass function (Stone & Metzger, 2016; Kochanek, 2016).
@@
 the Zwicky Transient Facility (ZTF) Graham et al., 2019; Bellm et al., 2019).
-They are slow (lasting 100100100100200200200200 days), blue (peaking in the UV) transients that do not typically cool as they fade (Gezari, 2021).
+They are slow (lasting 100100200200 days), blue (peaking in the UV) transients that do not typically cool as they fade (Gezari, 2021).
 TDEs last on scales of months, typically fading from peak in $\sim 200$ days (van Velzen et al., 2021).
@@
 They would indicate within which galaxies lower-mass SMBHs exist in the early universe, potentially distinguishing between SMBH seed models (Volonteri et al., 2021).
-Previously, Kochanek (2016) has predicted that high-redshift TDEs would not be observable in ground-based wide-field time domain surveys.
+Previously, Kochanek (2016) has used semi-empirical modeling of TDE rates to predict intrinsic and observed rates of TDEs as a function of redshift.
+They predict that, due to the observed “downsizing” of AGN mass over redshift (as modeled in Shankar et al., 2009), the intrinsic TDE rate will decrease by a factor of $5$ by $z=1$.
+As a result, they assert that the observed rate of high-redshift TDEs, given ground-based wide-field time domain surveys (focusing on Pan-STARRS1 Chambers et al., 2016), will be even lower.
 Here, we extend to high redshifts by leveraging deep space-based imaging without the use of time series.
@@
 (in prep.) that, given the observed evolution of galaxy and SMBH properties with redshift, the TDE rate at high redshifts is likely greatly enhanced.
-As a result, in upcoming surveys it may be possible to study populations of TDEs at high redshift.
+As a result, in upcoming surveys such as the Roman Space Telescope High Latitude Time Domain Survey (Rose et al., 2021) and the Rubin Observatory’s Legacy Survey of Space and Time (LSST Ivezić et al., 2019) it may be possible to study populations of TDEs at high redshifts.
 Inayoshi et al.
@@
 We present the properties of HZTDE-1, a transient high-redshift TDE candidate in Section 4.
-We compare potential supernova models for this object to a TDE explanation.
+We compare the TDE model to potential supernova models for this object.
 We discuss the need for spectroscopy of this object, and its implications if validated in Section 5.
 We summarize our work in Section 6.
-[…]
-3
 2 Identifying High-Redshift TDEs in Deep Surveys
@@
 To first order, the UV/optical SED of a TDE is well-approximated by a black body with a typical temperature of $\sim 10^{4.3\pm 0.15}~K$ (van Velzen et al., 2020, 2021; Hammerstein et al., 2023; Yao et al., 2023).
-We simulate TDEs with the the same temperature distribution as the latest sample of spectroscopically confirmed TDEs from ZTF (Yao et al., 2023), and simulate their volumetric density using the ZTF $g$-band luminosity function.
+We simulate TDEs with the the same temperature distribution as the latest sample of spectroscopically confirmed TDEs from ZTF (Yao et al., 2023), and simulate their volumetric density by integrating over the double-power law fit to the ZTF $g$-band luminosity function (their Equation 16).
 TDEs typically do not cool as they fade, so we assume constant-temperature blackbodies for their entire evolution.
@@
 $L_{\nu}(t)$ $=L_{\nu_{0}\,\rm peak}~\frac{B_{\nu}(T_{0})}{B_{\nu_{0}}(T_{0})}$
-×$\times\begin{cases}e^{-(t-t_{\rm peak})^{2}/2\sigma^{2}}&t\leq t_{\rm peak}\\ e^{-(t-t_{\rm peak})/\tau}&t>t_{\rm peak}\\ \end{cases}$ (1)
-with the luminosity in the $g$-band in erg s-1 $44.6842.68<\log(L_{g})<$ 44.68, the Gaussian rise time $1.30.4<\log(\sigma)<$ 1.3 days and the exponential decay time $2.31.2<\log(\tau)<$ 2.3 days from ZTF.
-We allow TDEs to fade until they reach the magnitude limit of the filter/survey observing them, to account for the time evolution of the TDE.
+[…]
+with the luminosity in the $g$-band in erg s-1 $42.68<\log(L_{g})<44.68$, the Gaussian rise time $0.4<\log(\sigma)<1.3$ days and the exponential decay time $1.2<\log(\tau)<2.3$ days from ZTF.
+When simulating TDEs, we step over redshift in shells of size $\Delta z=0.2$.
+At a given $z$, we sample over their temperatures and luminosities in a grid starting with the faintest TDE at a given redshift that is visible given the survey limits.
+We sample their luminosities in step sizes inversely proportional to their probability densities, giving more dense sampling to the higher-probability luminosities.
+At each redshift and luminosity, we simulate a TDE with every temperature observed in the Yao et al.
+(2023) ZTF TDE sample.
+At each given luminosity, we allow TDEs to fade as $L\propto t^{-5/3}$ until they reach the magnitude limit of the filter/survey observing them, to account for the time evolution of the TDE.
 2.1 The Host model
+Figure 1: Apparent magnitudes of the median TDE from our simulations (solid lines) compared to the median ZTF host galaxy (dashed lines Hammerstein et al., 2021).
+The shaded regions represent the 5th and 95th percentile for the TDE hosts.
+95% of the hosts drop out by $z\sim 2.2$ in F115W, and by $z\sim 4.1$ in F [^4] W.
 We base our predictions of properties of TDE host galaxies on local observations of TDE hosts.
@@
 (2021), which are estimated by maximum a posteriori fits of Flexible Stellar Population Synthesis models (Conroy et al., 2009) to broadband photometry from SDSS, Pan-STARRS, and GALEX.
-When calculating properties of TDE hosts, we simulate every known TDE host to account for their full diversity of properties.
-When simulating observed TDEs in a survey, we include the host spectrum by adding a mean template, derived from the Prospector models, to the TDE spectrum before performing synthetic photometry.
+When calculating properties of TDE hosts, we simulate every TDE host from ZTF to account for their full diversity of properties.
+We find that, due to the k-correction and the homogeneous colors of the TDE hosts, they typically are not visible beyond $z=3.5$ as seen in Figure 1.
 We also use these ZTF TDE hosts as a basis for the physical sizes for typical TDE hosts.
@@
 As NIRCam has a resolution of 0.07 arcsec at 2 microns, all TDE host galaxies above the survey magnitude limit will be resolvable as an extended source in COSMOS-Web.
-At early epochs ($z\sim 7$), galaxies are expected to be intrinsically smaller and fainter (Grazian et al., 2012; Kawamata et al., 2018), such that even TDEs in unusual hosts will likely be point sources.
-Because of this, we predict that the majority of TDEs at high redshifts will outshine their hosts in the COSMOS-Web filters and be point sources.
+At early epochs ($z\sim 7$), galaxies are expected to be intrinsically smaller (Grazian et al., 2012; Kawamata et al., 2018), such that even TDEs in unusually bright or blue hosts will not have spatially extended emission.
 Because all of the host galaxies we observe are at $z<1$, the galaxies we use are of higher mass and more luminous than those which we expect at high redshifts.
@@
 While there is a minority of these galaxies which are exceptionally luminous without SMBHs above the Hills mass due to extremely active star formation (e.g.
-Franco et al., 2024; Casey et al., 2024; Castellano et al., 2024), it would be extremely difficult to detect a TDE in one because its star formation is bright and blue, making the TDE emission a sub-dominant component of its UV SED.
+Franco et al., 2024; Casey et al., 2024; Castellano et al., 2024), it would be extremely difficult to detect a UV/optical TDE in one because its star formation is bright and blue, making the TDE emission a sub-dominant component of its UV SED.
 Therefore, we do not consider them in our analysis.
-This makes the local TDE hosts we observe effectively upper limits for higher-redshift host galaxy luminosity.
-In summary, TDEs trace lower-mass SMBHs which trace, in turn, lower mass galaxies.
-These are expected to be even less massive at high-redshifts, meaning we expect the majority of TDEs to be hostless.
+However, a population of TDEs which occur primarily in star-forming galaxies have been observed to be infrared-bright with no UV/optical component to their emission (Masterson et al., 2024), making their rates relevant to wide-field surveys using MIRI.
+For our search in COSMOS-Web, the local TDE hosts we observe are effectively upper limits for higher-redshift host galaxy luminosity.
+In summary, TDEs trace lower-mass SMBHs which trace, in turn, lower-mass galaxies.
+These are expected to be even less massive at high redshifts, therefore intrinsically fainter, and are k-corrected into the mid-infrared.
+We therefore expect the vast majority of TDEs at $z>4$ to appear hostless.
 2.2 Dust extinction
@@
 (mag) (mag) (mag) (mag)
-F115W - F [^4] W F150W - F277W $0.0024$ $0.4766$ $0.0473$ $-1.4$ 3.5 $-1.1$ $1.5$
-F115W - F150W F277W - F [^4] W $0.0242$ $0.5215$ $-0.6974$ $-0.4$ $1.3$ $-1.0$ $0.1$
-F115W - F150W F150W - F277W $0.0605$ $1.3533$ $-0.5377$ $-0.4$ $1.3$ $-1.1$ $1.5$
-F150W - F277W F277W - F [^4] W $0.0005$ $0.3859$ $-0.4900$ $-1.1$ $1.5$ $-1.0$ $0.1$
+F115W - F [^4] W F150W - F277W $0.0024$ $0.4766$ $0.0473$ 1.4-1.4 3.5 1.1-1.1 $1.5$
+F115W - F150W F277W - F [^4] W $0.0242$ $0.5215$ 0.6974-0.6974 0.4-0.4 $1.3$ 1.0-1.0 $0.1$
+F115W - F150W F150W - F277W $0.0605$ $1.3533$ 0.5377-0.5377 0.4-0.4 $1.3$ 1.1-1.1 $1.5$
+F150W - F277W F277W - F [^4] W $0.0005$ $0.3859$ 0.4900-0.4900 1.1-1.1 $1.5$ 1.0-1.0 $0.1$
 Table 1: Parabolas of best fit to the high-redshift TDE color-color relationship, as seen in Figure 4 anr parameterized in Eqquation 2.
@@
 We also fit bounds to each individual color, which, together with the parabola fits, are presented in Table 1.
-We fit a parabola to the color-color relationship of TDEs in each filter combination, parameterized as:
+We fit a parabola to the color-color relationship of TDEs in each filter combination through maximum likelihood optimization with the assumption of uniform uncertainties, which is mathematically equivalent to least squares fitting.
+Each curve is parameterized as:
 $C_{2}=a_{1}C_{1}^{2}+a_{2}C_{1}+a_{3}$ (2)
@@
 The values of coefficients $a_{1}$, $a_{2}$, and $a_{3}$ for each pair of colors are listed in Table 1, alongside the bounds on $C_{1}$ and $C_{2}$ form our simulations.
+We place bounds on the redshift range we simulate, on the low end by the redshift beyond which TDEs will be point sources, and on the high end by the redshift at which Lyman absorption renders them undetectable.
 Additionally, we fit the minimum apparent magnitude expected for a TDE at any redshift of a given color, shown in Figure 4, as
@@
 To check our blackbody idealization, we take the following TDEs that have Hubble UV spectroscopy, and simulate them at high redshifts: iPTF16fnl, ASASSN-15oi, AT2022dsb, AT2018bsi, and ASASSN-14li (Charalampopoulos et al., 2022; Cenko et al., 2016; Blagorodnova et al., 2019; Brown et al., 2018; Hung et al., 2019; Hoogendam et al., 2024).
-We extend the UV spectrum into the optical ($>2600$) with a blackbody with the temperature fit by Mummery & van Velzen (2024), redshift the TDE spectra, and add in Lyman absorption shortward of rest-frame 1216Å (Inoue et al., 2014).
+The spectra used to create these templates can be accessed via doi: 10.17909/9m1b-2s28.
+We extend the UV spectrum into the optical (>2600>2600 Å) with a blackbody with the temperature fit by Mummery & van Velzen (2024), redshift the TDE spectra, and add in Lyman absorption shortward of rest-frame 1216Å (Inoue et al., 2014).
 We show these templates in Figure 3.
@@
 The red line is the HST spectrum, purple dotted is the blackbody used to extend the spectrum, and photometry is in blue.
-We switch from the HST spectrum to the blackbody at rest-frame wavelength $2600$, which is shown in the changing opacity of each line.
+We switch from the HST spectrum to the blackbody at rest-frame wavelength $2600$ Å, which is shown in the changing opacity of each line.
 We simulate this spectrum at higher redshifts to test our color cuts.
 2.4 Alternative TDE Models
-By extrapolating the blackbody NUV/optical spectrum of a TDE to the UV, we predict the spectrum of TDEs at rest-frame wavelengths at which their spectra are relatively poorly constrained, between $900$ and $1500$.
+By extrapolating the blackbody NUV/optical spectrum of a TDE to the UV, we predict the spectrum of TDEs at rest-frame wavelengths at which their spectra are relatively poorly constrained, between $900$ Å and $1500$ Å.
 Furthermore, the presence of Bowen lines in TDE spectra (e.g.
-Leloudas et al., 2019) implies that the flux predicted by a blackbody at $228$ is $\sim 5$ orders of magnitude less than what is needed in order to ionize He II and cause Bowen fluorescence.
-In the intermediate wavelength that we observe, between $900$ and $1500$, a blackbody may also under-predict the TDE flux.
+Leloudas et al., 2019) implies that the flux predicted by a blackbody at $228$ Å is $\sim 5$ orders of magnitude less than what is needed in order to ionize He II and cause Bowen fluorescence.
+In the intermediate wavelength that we observe, between $900$ Å and $1500$ Å, a blackbody may also under-predict the TDE flux.
 To account for this potential discrepancy, we also consider a TDE spectrum with enough UV flux in order to trigger Bowen lines.
@@
 (2018) model, which uses 3D GRRMHD simulations to simulate a super-Eddington accretion disk, and Monte Carlo radiative transfer to simulate reprocessing; we do not claim that this model is correct but treat it as an example of a spectrum which is extremely luminous in the UV.
-We select emission from a viewing angle between 22.5 $22.5-\>45^{\circ}$ from a simulated spectrum which matches a blackbody spectrum of a few ×$\times 10^{4}$ K in the optical.
+We select emission from a viewing angle between $22.5-\>45^{\circ}$ from a simulated spectrum which matches a blackbody spectrum of a few $\times 10^{4}$ K in the optical.
 This model is not used for any rate calculations or final conclusions, but is a comparison point for TDE colors and identification methodology.
@@
 Lower redshift supernovae: SNe Ia, SNe IIP and SNe IIn at $z>2$ may be point sources in JWST data.
+Already, supernovae with host redshifts from 0.6461 to 3.6 have been discovered using NIRCam imaging in the JWST Advanced Deep Extragalactic Survey (JADES) (Pierel et al., 2024; Siebert et al., 2024; DeCoursey et al., 2024; Coulter et al., 2025).
 We specifically model SNe which are point sources, so we do not consider contribution in their colors from host galaxy light.
+Our SN Ia models are SALT3 models which extend into the NIR (Pierel et al., 2022), with parameter sweeps over $2\leq x_{1}\leq 2$, $0.3\leq c_{1}\leq 0.3$, and $z\leq 5$, simulated at phases $20\leq\textrm{p}\leq 50$ days.
 We predict that no SNe Ia will be detected beyond $z>5$ in COSMOS-Web.
-Already, supernovae with host redshifts from 0.6461 to 3.6 have been discovered using NIRCam imaging in the JWST Advanced Deep Extragalactic Survey (JADES) (Pierel et al., 2024; Siebert et al., 2024; DeCoursey et al., 2024; Coulter et al., 2025).
-To model these, we use from SNANA SNe II models (Kessler et al., 2009) of observed SNe at their peak luminosities, which have been implemented in sncosmo.
-Our SN Ia models are SALT3 models which extend into the NIR (Pierel et al., 2022), with parameter sweeps over $-2\leq x_{1}\leq 2$, $-0.3\leq c_{1}\leq 0.3$, and $z\leq 5$, simulated at phases $-20\leq\textrm{p}\leq 50$ days.
-The SNANA have only an amplitude parameters, so we sweep over a range of luminosities for each template.
-We set bounds on each supernova subtype’s luminosity distribution based on observed samples, $-16<M_{b}<-19$ for SNe Ia (Phillips, 1993), $-16\leq M_{V}\leq-19.5$ for SNe Ib/c (Drout et al., 2011), $-16\leq M_{b}\leq-18.5$ for SNe IIP (Sanders et al., 2015), $-13\leq M_{g}\leq-22$ for SNe IIn (Nyholm et al., 2020), and $-20\leq M_{g}\leq-22$ for SLSNe (Gomez et al., 2024).
-We find that SNe Ia are sufficiently bluer than TDEs at lower redshifts that match their apparent magnitudes, and there is no confusion at any redshift.
-SNe Ib/c can indeed match the colors and magnitude of high-redshift TDEs at redshifts $z<1$, however their hosts are expected to be detected in COSMOS-Web and previous deep optical imaging of the COSMOS field.
-SNe II are even bluer, and can pass as $z>4$ TDEs around $z\sim 2$, but would need rest-frame $b$-band absolute magnitudes $\lesssim-20$, which is not observed in SNe IIP.
+To model SNe II, we use every SNANA model (Kessler et al., 2009) that has been implemented in sncosmo to represent the diversity of observed SNe II at their peak luminosities.
+The SNANA models have only an amplitude parameters, so we sweep over a range of luminosities for each template.
+We set bounds on each supernova subtype’s luminosity distribution based on observed samples, $16<M_{b}<-19$ for SNe Ia (Phillips, 1993), $16\leq M_{V}\leq-19.5$ for SNe Ib/c (Drout et al., 2011), $16\leq M_{b}\leq-18.5$ for SNe IIP (Sanders et al., 2015), $13\leq M_{g}\leq-22$ for SNe IIn (Nyholm et al., 2020), and $20\leq M_{g}\leq-22$ for SLSNe (Gomez et al., 2024).
+We find that SNe Ia are sufficiently bluer than TDEs at lower redshifts where their apparent magnitudes match, and are too faint at higher redshifts when their colors match, such that there is no confusion at any redshift as seen in Figure 4.
+SNe Ib/c can indeed match the colors and magnitudes of high-redshift TDEs at redshifts $z<1$, however their hosts are expected to be detected in COSMOS-Web and previous deep optical imaging of the COSMOS field.
+SNe II are even bluer, and at $z\sim 2$ they can pass as $z>4$ TDEs, but would need rest-frame $b$-band absolute magnitudes $\lesssim-20$, which is not observed in SNe IIP.
 SNe IIn can become this luminous at maximum brightness, serving as our main contaminant if their hosts are not detected at $z\sim 2$.
+We show SNe IIn at $1.5<z<2.5$ in Figure 4.
 Superluminous supernovae (SLSNe): SLSNe, alongside TDEs and pair-instability supernovae, will be among the few transients which are visible up to very high redshifts.
 We model them using the templates constructed from the Dark Energy Survey (DES) SLSNe, which have a blackbody continuum with UV absorption features superimposed (Angus et al., 2019).
-Because their absorption strengths evolve and the SLSNe cool after their explosion, we simulate them up from $-10$ to $+35$ rest-frame days from peak with early-time temperatures between $5000$ K and $18000$ K.
+Because their absorption strengths evolve and the SLSNe cool after their explosion, we simulate them up from 10-10 to +35+35 rest-frame days from peak with early-time temperatures between $5000$ K and $18000$ K.
 We use sncosmo to model these and perform synthetic photometry (Barbary et al., 2016).
@@
 Time domain information, or spectroscopy is needed to distinguish a candidate TDE from a SLSN: SLSNe typically cool as they fade, whereas TDEs remain constant in color.
-A fading source with no color changes in two to three total observations over rest-frame months (observer-frame 66668888 months between observations) is an extremely strong TDE candidate.
+A fading source with no color changes in two to three total observations over rest-frame months (observer-frame 6688 months between observations) is an extremely strong TDE candidate.
 Finally, spectroscopy can identify typical TDE emission lines, such as broad H-alpha and He II, which can indicate a TDE (Gezari, 2021).
@@
 We calibrate the absolute magnitude of the model in the $r$-band to the high-redshift ($z=7$) QSO luminosity function in Matsuoka et al.
-(2023), with a break magnitude of $-25.6$ mag.
+(2023), with a break magnitude of 25.6-25.6 mag.
 We redshift this quasar model at redshifts up to z=8, just beyond the highest-redshift confirmed quasar, J0313–1806 (Wang et al., 2021).
@@
 Neither serves as a major contaminant in our sample.
-Stars: We implement the PHOENIX stellar templates for stars with $5.50<\log{g}<$ 5.5, $240002000<T_{\textrm{eff}}<$ 24000 and $-2<\textrm{Fe}/\textrm{H}<1$ (Husser et al., 2013).
+Stars: We implement the PHOENIX stellar templates for stars with $0<\log{g}<5.5$, $2000<T_{\textrm{eff}}<24000$ and $2<\textrm{Fe}/\textrm{H}<1$ (Husser et al., 2013).
 We use pysynphot both to generate their spectra and to perform synthetic photometry in the NIRCam filters.
 We later use the same PHOENIX models to carefully simulate cool dwarf stars, such as M and L dwarfs.
-We simulate cool dwarfs to Milky Way halo distances, up to $20$ kpc, although distances $>5$ kpc are unlikely due to the scale height of the Milky Way, or until they are too faint for detection in COSMOS-Web.
+We simulate cool dwarfs to Milky Way halo distances, up to $20$ kpc, although distances >5>5 kpc are unlikely due to the scale height of the Milky Way, or until they are too faint for detection in COSMOS-Web.
 Ultracool dwarf stars with temperatures $2000\leq T_{BB}\leq 3200$ K can masquerade as high redshift TDEs because they have predominantly thermal cool continuum emission which resembles redshifted hot thermal emission; e.g.
 a 2400 K local blackbody has the same spectral shape as a $z\sim 6$ 15000 K blackbody.
-While the dwarf star spectral templates we have are slightly flatter than blackbodies, they become a contaminant when searching for higher-redshift ($z>6$), cooler ($<20000$ K) TDEs.
+While the dwarf star spectral templates we have are slightly flatter than blackbodies, they become a contaminant when searching for higher-redshift ($z>6$), cooler (<20000<20000 K) TDEs.
 To distinguish these, we need to detect either time evolution, or a Lyman break to verify a candidate as an extragalactic transient.
-This is most prevalent when selecting bright ($>25$ mag) sources, but a halo star could hypothetically mimic a faint TDE candidate.
-Brown dwarfs: We use the Sonora Bobcat models (Marley et al., 2021). These are both rare, and they do not match colors of TDEs so this comparison is just for verification of their difference.
+This is most prevalent when selecting bright (>25>25 mag) sources, but a halo star could hypothetically mimic a faint TDE candidate.
+Brown dwarfs: We use the Sonora Bobcat models (Marley et al., 2021).
+These are both rare, and they do not match colors of TDEs so this comparison is just for verification of their difference.
+They are omitted from Figure 4 for visual clarity, because they lie far outside the axis limits.
 Point-like galaxies: A small fraction of very-high-redshift galaxies observed in JWST are compact enough to render them point-like.
@@
 This catalog models the evolving luminosity functions, spectroscopic properties, and morphologies of galaxies up to $z\sim 15$, and can reproduce observations up to $z\sim 10$.
-We specifically use simulations which use templates of expected $155<z<$ 15 galaxies.
+We specifically use simulations which use templates of expected $5<z<15$ galaxies.
 We select galaxies which will be point sources by selecting those which have an observed angular size equal to or less than the FWHM of the F115W PSF, $0.07$ arcsec.
 We use the pre-synthesized photometry from these catalogs for a wide range of galaxy types and redshifts, up to $z=15$.
-We find that $>99\%$ of unresolved galaxies simulated will be too faint, given their color, to resemble a TDE as seen in Figure 3.
+We find that $99\%$ of unresolved galaxies simulated will be too faint, given their color, to resemble a TDE as seen in Figure 4.
 However, rare unresolved extremely UV-bright galaxies are possible (e.g.
@@
 HZTDE-1 is notably brighter than these templates because the majority of the template UV spectra are taken weeks to months after maximum brightness.
+The (Dai et al., 2018) model is represented in the red-blue line, with the bluest end representing $z=2$, and the reddest representing $z=8$.
 To summarize, our main persistent contaminants are very bright compact blue point source galaxies, local M- and T-dwarf stars, and hostless superluminous supernovae.
@@
 To rule out a superluminous supernova photometrically a second epoch of imaging is needed to observe the decline of the source.
-[…]
-3
 3 COSMOS-Web Search Methodology
@@
 The COSMOS-Web survey images $0.54$ deg2 of the sky with NIRCam to depths of $27.45$ mag in F115W, $27.66$ mag in F150W, $28.28$ mag in F277W and $28.17$ mag in F [^4] W, and occasionally in the Mid-Infrared Instrument (MIRI) band F770W to $\sim 26$ mag (Casey et al., 2022).
-Decades of deep multiwavelength observations of the COSMOS field means that, for a given source, one can look back over 20 years of observations (e.g.
+Decades of deep multi-wavelength observations of the COSMOS field means that, for a given source, one can look back over 20 years of observations (e.g.
 Scoville et al., 2007; Capak et al., 2007; Koekemoer et al., 2011).
@@
 (2023), this rate could be much higher.
-We apply our method to COSMOS-Web catalog (M.
-Shuntov, L.
-Paquereau et al.
-in prep.; see Akins et al.
-(2024) for details on catalog construction): we simulate TDEs in the NIRCam filters F115W, F150W, F277W and F [^4] W using the methods described in Section 2, apply cuts based on their colors, magnitudes, and morphologies, and then check in previous observations to verify if these sources are indeed transient.
+We apply our method to COSMOS-Web catalog (Shuntov et al., 2025), which combines new and archival imaging of the COSMOS field.
+Details on NIRCam image reduction can be found in Franco et al.
+(2025).
 To select point sources, we calculate the ratio between the flux in a $0.5$ arcsecond aperture and in a $0.25$ arcsecond aperture.
@@
 We restrict only to objects which are point-like in F115W and F150W, the filters in which we most confidently expect no galaxy light beyond $z\sim 4$.
-Figure 4: Flux ratio between a $0.25"$ aperture and $0.5"$ aperture in F150W vs magnitude for all sources in the COSMOS-Web catalog (gray points).
+Figure 5: Flux ratio between a 0.250.25" aperture and 0.50.5" aperture in F150W vs magnitude for all sources in the COSMOS-Web catalog (gray points).
 Sources which are classified as stars via LePHARE SED fits (blue points) are used as a point source selection guide.
@@
 (2022) Table 1), and is publicly archived.
-In this analysis, we use forced photometry for previous surveys at the location of a detection in COSMOS-Web, and classify a nondetection as any source with a $<3\sigma$ detection in flux.
+In this analysis, we use forced photometry for previous surveys at the location of a detection in COSMOS-Web, and classify a nondetection as any source with a $3\sigma$ detection in flux.
+Details on the forced photometry performed for all COSMOS-Web sources can be found in the COSMOS-Web catalog paper (Shuntov et al., 2025).
 This is useful when a TDE candidate is sufficiently bright such that a non-detection in shallower ground-based data is constraining; the detection image in COSMOS2020 is a $izYJHK_{s}$ stack.
 We use these data to extend SEDs when visualizing candidates, but do not use them in our candidate selection or fitting in case of variability between the HSC observations and COSMOS-Web.
-We search for our candidates from COSMOS-Web in the COSMOS2020 catalog using the IPAC COSMOS Archive (COSMOS Project, 2020) for detections in existing COSMOS catalogs within a matching radius of 1”.
+We search for our candidates from COSMOS-Web in the COSMOS2020 catalog using the IPAC COSMOS Archive (doi: 10.26131/irsa178)for detections in existing COSMOS catalogs within a matching radius of 1”.
 For the subset of candidates which are not in COSMOS2020, we download their cutouts and manually inspect them.
@@
 We ultimately selected 1764 sources that passed our point source cuts, 117 passed our color-color, and magnitude cuts.
+We select sources based on COSMOS-Web NIRCam colors alone, in order to prevent time evolution between different surveys from affecting calculated colors.
 We visualize these candidates in color-color and color-magnitude space in Figure 6.
@@
 “Bright” sources are sources with $M_{F115W}>27.8$ AB mag, where $27.8$ mag is the detection limit in HST F814W in COSMOS2020, the deepest imaging of this field.
-These filters do not overlap– the $\lambda_{\rm max}$ of F814W is $9643$ and the $\lambda_{\rm min}$ of F115W is $9975$– but they are near enough to where nondetection in F115W approximates a nondetection in F814W.
+These filters do not overlap– the $\lambda_{\rm max}$ of F814W is $9643$ Å and the $\lambda_{\rm min}$ of F115W is $9975$ Å– but they are near enough to where nondetection in F115W approximates a nondetection in F814W.
 Therefore, “bright” sources should have mostly been detected in HST.
@@
 For all 79 sources without HST catalog entries, we download the HST F814W cutouts, and search through them by hand for nondetections.
-We ultimately determine by visual inspection that 72 of the 79 “bright” sources which pass our cuts do have a counterpart in F814W images, but are excluded from COSMOS2020.
+We ultimately determine by visual inspection that 72 of the 79 “bright” sources which pass our cuts do indeed have a counterpart in F814W images, but are excluded from COSMOS2020.
 We note that these sources often lie on the edge of a tile, in the masked region near a bright star, or have inaccurate astrometry.
-Among the sources not seen in the manual inspection, six are very red and not bright in F184W.
+Among the sources not seen in F814W in manual inspection, six of them appear to be static, yet are too faint in F814W to be detected by previous HST imaging.
 We are left with one source which is detected in COSMOS-Web but not in COSMOS2020.
-The remaining candidate peaks at $\sim 25$ mag in F150W, which is $>1$ mag above the limit of previous infrared imaging with UltraVISTA.
+The remaining candidate peaks at $\sim 25$ mag in F150W, which is >1>1 mag above the limit of previous infrared imaging with UltraVISTA.
 This source is not detected in UVISTA data to even $1\sigma$, confirming its nature as a transient and viable high-redshift TDE candidate.
@@
 Despite having a blackbody-like SED shape in NIRCam, the HSC photometry shows that the SED is flat in the optical, with a Lyman break beginning in the g-band.
-Its F770W fluc greatly exceeds that of a blackbody.
+Its F770W flux greatly exceeds that of a blackbody.
 The source appears point-like in F115W and F150W, but in the F277W imaging it can be seen to be slightly extended, although compact.
 The best-fit galaxy template is plotted in red, star template in light blue, and quasar in purple.
+This galaxy does not have a F814W flux measurement in the COSMOS2020 catalog due to photometry flags, but manual inspection confirms it was detected.
 Bottom: Example of a high-redshift TDE candidate which we believe to be a local cool dwarf star.
@@
 This source has no Lyman break, and is detected to $5\sigma$ in every HSC band.
-These candidates are selected only by JWST photometry, and the HSC photometry is consistent with a blackbody with the same peak as a high redshift TDE for the vast majority of sources, as seen in Figures 6 and 7.
+These candidates are selected only by JWST photometry, and the HSC photometry is consistent with a blackbody with the same peak as a high redshift TDE for the vast majority of sources, as seen in Figures 7 and 8.
 Figure 8 shows an example fit of our simplified high-redshift TDE model to HZTDE-1, characterized only by absolute magnitude, blackbody temperature, and redshift, to the photometry.
@@
 This is expected for a high-redshift TDE, because their soft X-ray spectrum would be completely absorbed by the CGM beyond $z\sim 3$.
-Objects without Lyman dropout are likely cool dwarf stars.
-Stars with temperatures $2400\textrm{K}<T<3500\textrm{K}$ have blackbody peaks around 1.2 $\mu$m, the same observed wavelength as a $z>5$ TDE.
-These stars are abundant in the Milky Way, although are less common when looking out of the Galactic Plane.
-Figure 6 shows a source which we believe is a star.
-It has no listing in the COSMOS2020 catalog, but manual inspection can identify the corresponding object in previous COSMOS imaging.
+[…]
 4 Properties of the TDE Candidate HZTDE-1
@@
 The cutouts include imaging from UltraVISTA, Hyper Suprime Cam, Hubble ACS, and JWST NIRCam.
-The source is only detected to $>3\sigma$ in the NIRCam images.
-The SED shows the observed SED as red circles, synthetic photometry for the best-fit galaxy spectrum (purple line) and TDE spectrum (red line) in diamonds, and all photometric upper limits as downward-pointing triangles.
+The source is only detected to $3\sigma$ in the NIRCam images.
+The SED shows the observed SED as red circles, the best fit SLSN spectrum (blue line) and TDE spectrum (red line), and all photometric upper limits as downward-pointing triangles.
 Errorbars in magnitude exist for the NIRCam data but are too small to be seen.
 The transmission curve for each filter is plotted below it.
-A simulated spectrum of TDE 2019qiz at $z=5$ is overplotted as a dotted line, and a template for a $24,000$ K superluminous supernova at $z=3.6$ is the yellow line.
+A simulated spectrum of TDE 2019qiz at $z=4$ is overplotted as a dotted line.
 The TDE candidate HZTDE-1 is shown in Figure 8.
@@
 There are three nearby galaxies that HZTDE-1 could belong to if it is a lower-redshift supernova.
-We assess its possible association with the galaxy directly above it in the cutouts shown in Figure 7, because it is the closest galaxy, with an angular separation of $0.998$”, the brightest galaxy with $M_{F444W}=23.7$ AB mag, and the most extended with effective radius $R_{e}=0.20$” as modeled by GALFIT (Peng et al., 2002).
+We assess its possible association with the galaxy directly above it in the cutouts shown in Figure 8, because it is the closest galaxy, with an angular separation of 0.9980.998”, the brightest galaxy with $M_{F444W}=23.7$ AB mag, and the most extended with effective radius $R_{e}=0.20$” as modeled by GALFIT (Peng et al., 2002).
 The GALFIT model for the sources in the field is shown in Figure 10, alongside the F [^4] W image and the residual.
 The best-fit galaxy spectrum to this source’s SED has a photometric redshift $z_{\rm phot}=1.75^{+0.1}_{-0.07}$.
+This galaxy has a slightly above average best-fit specific star formation rate given its photometric redshift, of $\log(\rm{sSFR})/M_{\odot}~yr^{-1}=-8.67$.
+An especially high star formation rate would support the occurrence of a supernova in its outskirts (Botticella et al., 2012).
 The SED is well-sampled from the HSC $g$-band to NIRCam F [^4] W, and the model spectrum fits accurately with a $\chi^{2}=1.72$ per degree of freedom.
@@
 They then calculate the $d_{\rm DLR}$, which is the distance between the supernova and the galaxy normalized by the DLR.
-They utilize both $HST$ ACS data and MICECATv2.0 simulations to show that any SN with a galaxy $>5$ DLR away should be assigned “hostless”, and achieve $\sim 95\%$ completeness with a similar cut.
+They utilize both $HST$ ACS data and MICECATv2.0 simulations to show that any SN with a galaxy >5>5 DLR away should be assigned “hostless”, and achieve $\sim 95\%$ completeness with a similar cut.
 We calculate the unitless dDLR for HZTDE-1 to be $5.05$ relative to the nearest galaxy, making it highly likely to be a hostless transient.
-This dDLR is even higher for the short-wavelength NIRCam filters ($>6$ DLR in F115W), which have higher spatial resolution.
+This dDLR is even higher for the short-wavelength NIRCam filters (>6>6 DLR in F115W), which have higher spatial resolution.
 Figure 10: GALFIT modeling of the morphology of the sources in the F [^4] W imaging of HZTDE-1.
-We display the size of $1$ directional light radius (DLR) from the galaxy to the transient to show the accuracy of the modeling.
-The HZTDE-1 is $>5$ DLR from the galaxy, unlikely to be associated.
+We display the size of 11 directional light radius (DLR) from the galaxy to the transient to show the accuracy of the modeling.
+The HZTDE-1 is >5>5 DLR from the galaxy, unlikely to be associated.
 In shorter-wavelength filters, this DLR separation increases.
-We assume that HZTDE-1 is unlikely to be a “hostless” non-superluminous supernova.
-However, given the rare case that HZTDE-1 is associated with the nearest galaxy despite extreme distance, we fit supernova SEDs for each supernova subtype at its photometric redshift, $z=1.75\pm 0.10$.
-We use the full range of supernova properties described in Section 2.5.
-For every supernova subtype, we fit every available model in sncosmo (Barbary et al., 2016), extending to the infrared with a blackbody when the model has insufficient wavelength range.
-We do not fit models for Population III SNe, Pair-Instability Supernovae (PISNe), or other exotic transients.
-We show our best-fit supernova spectra in Figure 10, and their respective lightcurves in Figure 11.
-We find that HZTDE-1 is significantly too red to be a SN Ia at $z\sim 1.75$ and too bright to be a SN Ia at $z>2$.
-HZTDE-1 is $\gtrapprox 1$ too bright to be explained by a Ib/c at $z\sim 1.75$.
-The SED is somewhat plausibly explained by an anomalously bright IIP before maximum brightness or a bright IIn prior to maximum brightness, but both under-predict the flux in F [^4] W at $z\sim 1.75$, which would be $1.6~{}\mu$m in the rest-frame.
-These best models are shown in Figure 10.
-This excess flux could be explained by contribution from thermal dust emission, but given the extreme offset from the potential host it is unlikely that the supernova would be embedded in dust.
-In summary, for HZTDE-1 to be a supernova associated with its nearest galaxy it must be an anomalously bright SN IIn at an extreme offset with an infrared flux excess.
+[…]
 Figure 11: Modeled spectra and synthetic photometry of best-fit Type Ia, Type Ib/c, IIP, and IIn supernovae compared to the SED of HZTDE-1.
@@
 They calibrate to the SLSN low-redshift rate, and account for the initial mass function and the fraction of massive stars to find that the SLSN volumetric rate increases with redshift until $z\sim 5$, where the SFR density peaks, and then the rate decreases with redshift indefinitely.
-At its maximum, $z\sim 5$, the volumetric SLSN rate is 10×10$\times$10 × the local rate.
-We expect the TDE rate to also be elevated by at least 10×10$\times$10 × the local rate, as shown in Karmen & et al.
+At its maximum, $z\sim 5$, the volumetric SLSN rate is $10\times$ the local rate.
+We expect the TDE rate to also be elevated by at least $10\times$ the local rate, as shown in Karmen & et al.
 (in prep.).
-HZTDE-1 would need to be extremely blue for a SLSN, with a best-fit temperature $24,000$ K at best-fit redshift $z=3.6$, but this is possible briefly before peak and during peak brightness.
-While this makes a SLSN somewhat unlikely, given the comparable rates there is no way to rule out a SLSN without further photometric monitoring or spectroscopy.
-Figure 11: Lightcurves of best-fit supernovae models to HZTDE-1, given that they are at redshift within $0.1$ of the closest galaxy.
-Each panel is in a single COSMOS-Web NIRCam filter.
-The best-fitting models, IIP and IIn, do not adequately explain the F [^4] W flux, which would be at $\sim 1.6\mu$m in the rest frame.
+HZTDE-1 would need to be blue for a SLSN, with a best-fit temperature 24, 000 K at best-fit redshift z = 3.6, but this is possible briefly before peak and during peak brightness.
+However, this fit was performed without a prior on SLSN behavior and a more reasonable 19,000 K SN will fit at z=3.2.
+Indeed, because of the low-mass nature of SLSN hosts, the SN could occur in a lower redshift, very faint host galaxy and would be intrinsically redder.
+Given the sample of host galaxies for the Palomar Transient Factory observed SLSNe (Perley et al., 2016), these hosts could drop out of the COSMOS-Web survey as soon as $z\sim 1.7$.
+This is around the peak of cosmic star formation rates and would imply a reasonable number of hostless SLSNe.
+Therefore, without further monitoring, the possibility of a lower-z SLSN is equally likely to a TDE.
+[…]
 5 Discussion and Future Confirmation
 In this work, we model the properties of TDEs at high redshift assuming that their populations are similar to those at low redshift.
-We argue that the assumptions made about their host galaxies (Section 2.1) and their colors (Section 2.3) are treated as bounds rather than constraints and that evolution of these properties with redshift will lie within our selection criteria.
+We argue that the assumptions made about their host galaxies (Section 2.1) and TDE spectra (Section 2.3) set limits on the colors and morphologies of high-redshift TDEs.
 However, any number of these assumptions can be relaxed.
 In future work, for example, modeling of TDEs in extended hosts, or modeling of dust-extincted TDE emission can increase the efficiency of searches for TDEs in wide-field surveys.
-As this search offers an opportunity to study the UV emission shortward of $2000$ which is usually not observed, deviations from blackbody assumptions about TDE spectra, such as those mentioned in Section 2.4 may emerge from studies of high-redshift TDEs.
-Both spectroscopic and photometric follow-up is needed to confirm HZTDE-1’s classification as a TDE and its redshift.
+As this search offers an opportunity to study the UV emission shortward of $2000$ Å which is usually not observed, deviations from blackbody assumptions about TDE spectra, such as those mentioned in Section 2.4 may emerge from studies of high-redshift TDEs.
+Either spectroscopic or photometric follow-up could be used to confirm HZTDE-1’s classification as a TDE and its redshift.
 TDEs have blue spectra with blackbody continua and broad Hydrogen and/or Helium emission features that narrow as the TDE evolves.
-A key distinction between TDEs and supernovae is their decline: TDEs fade slowly, typically as $t^{-5/3}$, and do not change temperature as they cool (van Velzen et al., 2020).
-Without imaging of a host galaxy to verify that HZTDE-1 is nuclear, combining identification these spectroscopic features and photometric monitoring is needed to confidently confirm it as a TDE.
-Spectroscopy, specifically in the infrared with JWST’s NIRSpec, is key to observing key TDE emission lines.
+A key distinction between TDEs and supernovae is their decline: TDEs do not change temperature as they cool (van Velzen et al., 2020), while supernovae cool rapidly after peak, typically 5000-8000 K (Gomez et al., 2024; Aamer et al., 2025).
+Without imaging of a host galaxy to verify that HZTDE-1 is nuclear, either identification of spectroscopic features or photometric monitoring is needed to confidently confirm it as a TDE.
+However, because of the degeneracy between redshift, luminosity, and temperature, a spectrum of the TDE or host would be needed to calculate a redshift.
+Spectroscopy, specifically in the infrared with JWST’s NIRSpec, is key to observing common TDE emission lines.
 X-ray emission is not expected because, as mentioned, the CGM will absorb all of the soft X-ray emission characteristic of TDEs at $z>3$.
 Longer wavelengths may be worth monitoring for later-time infrared flares due to dust echoes (van Velzen et al., 2016).
+If HZTDE-1 is a high-redshift SLSN, its discovery and classification will be valuable for understanding stellar evolution at higher redshift.
+SLSNe are expected to result from the explosions of massive stars (up to tens of solar masses Blanchard et al., 2020) and have a strong preference for galaxies with low metallicities (Perley et al., 2016; Schulze et al., 2018; Cleland et al., 2023).
+To first order, the cosmic rates of SLSNe should follow the cosmic star formation rate (Tanaka et al., 2013).
+Deviations from this relationship could provide insight into the abundance of metal-poor galaxies at earlier cosmic times.
+The methodology outlined in this paper could identify potentially hundreds of high-redshift SLSNe in upcoming wide-field surveys (e.g.
+the Roman High Latitude Wide Area Survey) and therefore could lead to measurements of how their rates and hosts change with redshift.
+This in turn can be used to study how massive stars form and evolve as the metallicity of the universe changes over cosmic time.
 5.1 TDE rates in COSMOS-Web
-In Karmen & et al.
+In the companion work Karmen & et al.
 (in prep.), we calculate observed TDE rates as a function of redshift in LSST, Roman, and the COSMOS-Web survey.
-We find that if the TDE rate does not evolve as a function of redshift, there is a $68\%$ chance of observing a single TDE in COSMOS-Web, and a $7\%$ chance of observing a TDE at $z>4$.
-We show, however, that the observed properties of high-redshift galaxies would imply a dramatically enhanced rate of TDEs at high-redshift $105<z<$ 10.
+We find that if we simply extrapolate the local TDE rate, scaled by the number density of SMBHs (Shankar et al., 2009) that can disrupt a main-sequence star (following Kochanek, 2016; Sun et al., 2015; Donnarumma et al., 2015), we get a $20\%$ chance of finding a single TDE in the COSMOS-Web survey.
+We show, however, that the observed properties of high-redshift galaxies would imply a dramatically enhanced rate of TDEs at high-redshift $z>5$, and perhaps as low as $z=2$.
 Namely, many galaxies are observed to be centrally compact (Finkelstein et al., 2023; Kartaltepe et al., 2023; Baggen et al., 2023; Guia et al., 2024), and may host nuclear stellar clusters (Ricotti et al., 2016; Sun et al., 2023) at high redshifts.
-This increases central stellar density, which in turn will lead to a ×$\times 100$ TDE rate enhancement (Kritos et al., 2024; Pfister et al., 2020).
+This increases central stellar density, $\rho_{\star}$, by 121-2 orders of magnitude (Baggen et al., 2024) which in turn can easily lead to a $\times 100$ TDE rate enhancement assuming a galaxy’s TDE rate scales as $\propto\rho_{\star}^{0.65}$ (Pfister et al., 2020), or even higher given the more dramatic observed scaling with stellar density $\propto\rho_{\star}^{0.9}$ (Graur et al., 2018).
+This enhancement is in agreement with simulations of TDEs in galaxies with nuclear stellar clusters (Kritos et al., 2024; Pfister et al., 2020).
 Additionally, galaxy mergers are shown to enhance TDE rates both observationally and theoretically.
@@
 With a galaxy merger rate that increases as a function of redshift, peaking at cosmic noon (Ventou et al., 2017; Duncan et al., 2019; Perna et al., 2025) we predict TDE rates to be enhanced from mergers as well.
-In the past it was expected that TDE rates decrease with the SMBH mass function as a function of redshift.
+In previous works (e.g.
+Kochanek, 2016; Donnarumma et al., 2015), the volumetric TDE rate has been scaled with a quickly-evolving SMBH mass function, as as a result it was expected that TDE rates decrease rapidly with the SMBH mass function as a function of redshift.
 However, given the abundance of supermassive and overmassive black holes detected at early cosmic times (e.g.
@@
 This methodology can be applied to any deep field survey, and will yield a wealth of high-redshift TDEs from the Roman High Latitude Wide Area Survey.
-  • We apply this selection method to the COSMOS-Web catalog, and identify 117 sources among $>700,000$ that pass our cuts.
+This methodology could also be used to identify a range of other hostless transients in upcoming surveys.
+  • We apply this selection method to the COSMOS-Web catalog, and identify 117 sources among >700,000>700,000 that pass our cuts.
 We systematically check these sources for lack of variability using the COSMOS2020 catalog, which compiles all previous imaging of the COSMOS field, and find one source which is a confirmed transient.
@@
   • We model HZTDE-1 as a possible supernova, both through SED modeling and host galaxy association.
-We find that HZTDE-1 is $>5$ directional light radii from the nearest galaxy, making an association extremely unlikely.
+We find that HZTDE-1 is >5>5 directional light radii from the nearest galaxy, making an association extremely unlikely.
 We find that while the SED cannot be explained by a Type I supernova, a Type II (particularly IIn) at $z\sim 2$ can explain the SED mildly well, except for the farthest infrared emission.
@@
   • We model HZTDE-1 as a possible TDE, and find that the most likely TDE has an effective blackbody temperature of $\log(T_{\rm{BB}})=4.31\pm 0.09$, absolute g-band magnitude of $M_{g}=-21.15^{+0.21}_{-0.13}$ and redshift of $z=5.02^{+1.33}_{-1.11}$.
-Follow-up observations of HZTDE-1, both for photometric fading and for spectroscopic information, are crucial for understanding its nature and confirming its redshift.
+Follow-up observations of HZTDE-1, for photometric fading and for spectroscopic information, are crucial for understanding its nature and confirming its redshift.
 If HZTDE-1 is a TDE, it could confirm an enhancement of the volumetric rate of TDEs in the high-redshift universe.
@@
 We then perform the same procedure using JWST imaging, with apertures of radius 3 pixels, as seen in Figure 13.
-As expected, we find a ×$\times 1000$ increase in flux relative to background in NIRCam, as seen in Figure 14.
+As expected, we find a $\times 1000$ increase in flux relative to background in NIRCam, as seen in Figure 15.
 Figure 13: Left: UltraVISTA $K_{s}$-band imaging of galaxy near HZTDE-1, surrounded by the apertures described in Appendix A.